### v3

# Approximation in Hölder spaces

Abstract.
We introduce new vanishing subspaces of the homogeneous Hölder space $\dot{C}^{0,\omega}(X,Y),$ in the generality of a doubling modulus $\omega$ and normed spaces $X$ and $Y.$ For many couples $X,Y,$ we show these vanishing subspaces to completely characterize those Hölder functions that admit approximations, in the Hölder seminorm, by smooth, Lipschitz and boundedly supported functions. Beyond the intrinsic interest of these results, we also present connections to bi-parameter harmonic analysis on the Euclidean space and in particular to compactness of the bi-commutator.
Key words and phrases:
smooth approximation, bounded support approximation, vanishing oscillation, Hölder space, Banach space
2020 Mathematics Subject Classification:
46T20, 41A29, 42B35, 46B20, 26B35
1. Introduction and main results
1.1. Introduction
We consider normed spaces $X,Y$ and moduli of continuity $\omega$ and study the Hölder spaces $\dot{C}^{0,\omega}(X,Y)$ that consist of those functions $f:X\to Y$ for which
$\|f\|_{\dot{C}^{0,\omega}(X,Y)}=\sup_{\begin{subarray}{c}x,y\in X\\ x\neq y\end{subarray}}\frac{\|f(x)-f(y)\|_{Y}}{\omega(\|x-y\|_{X})}<\infty.$
We establish various full characterizations of approximability in the class $\dot{C}^{0,\omega}(X,Y)$ by Lipschitz, smooth and boundedly supported functions in terms of three vanishing subspaces of $\dot{C}^{0,\omega}(X,Y)$ that demand vanishing behaviour of the Hölder norm of functions on small, large and far away scales. We will define these subspaces soon, but before that we make a few general observations.
First of all, if $Y$ is a Banach space, then so is $\dot{C}^{0,\omega}(X,Y),$ when equipped with the norm
$|||f|||_{C^{0,\omega}(X,Y)}=\|f\|_{\dot{C}^{0,\omega}(X,Y)}+\|f(0)\|_{Y}.$ (1)
Dropping the term $\|f(0)\|_{Y},$ the seminorm $\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)}$ becomes a norm when the functions are identified with equivalence classes modulo additive constants. In the particular case $\omega(t)=t^{\alpha},$ $\alpha\in(0,1),$ we recover the Hölder spaces $\dot{C}^{0,\alpha}(X,Y).$ Moreover, provided $\omega$ is any reasonable continuity moduli – at least $\omega(t)\to 0,$ as $t\to 0$ – the spaces $\dot{C}^{0,\omega}(X,Y)$ become subspaces of uniformly continuous functions.
The theory of uniform approximation of continuous functions by smooth functions has been extensively studied and developed in the last decades. For instance, when $X$ is a Hilbert space, Lasry and Lions [LasryLions] showed that Lipschitz real-valued functions can be uniformly approximated by $C^{1,1}$ Lipschitz functions. If $X$ is moreover separable, then a theorem of Moulis [Moulis] allows these approximations to be upgraded to $C^{\infty}$ approximations. Later, Cepedello-Boiso [Cepedello] obtained several results concerning $C^{1}$ or $C^{1,\alpha}$ approximation of uniformly continuous functions in super-reflexive Banach spaces $X$ – see also our Theorem 3.10 below and the very recent work of Johanis [Jo24], that contribute to this line of research. Then, Hájek and Johanis [HajJoh10] proved that when $X$ is separable with a $C^{k}$ smooth and Lipschitz bump, and either $X$ or $Y$ is super-reflexive, then uniformly continuous mappings admit approximations by $C^{k}$ maps that are still uniformly continuous. Lastly, we mention the real-analytic approximations of Azagra, Fry, and Keener [AzFrKe12] when $X$ is a Banach space with a separating polynomial; then, Lipschitz real-valued functions on $X$ can be uniformly approximated by real analytic Lipschitz functions. For more background references on smoothness and renorming on Banach spaces, and the related approximation results, we refer to the monographs [BenyaminiLindenstrauss], by Benyamini and Lindenstrauss; [DevilleGodefroyZizler], by Deville, Godefroy, and Zizler; and [HajJoh10], by Hájek and Johanis.
In the present article, we prove full characterizations of approximation with Hölder type rate of convergence taking the place of uniform convergence. We note that these two types of convergence do not follow from each other; see Appendix A for a detailed comparison with simple observations and examples. Our main result is best described as the recognition of the vanishing scales small, large and far (below) as precisely the concept that allows a complete description of the closure of smooth, Lipschitz and boundedly supported functions with respect to the $\dot{C}^{0,\omega}$ seminorm. We notate
$\operatorname{osc}^{\omega}_{\delta}(f)=\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|_{X}=\delta\end{subarray}}\frac{\|f(x)-f(y)\|_{Y}}{\omega(\|x-y\|_{X})},\qquad\operatorname{osc}^{\omega}_{(x,y)}(f)=\frac{\|f(x)-f(y)\|_{Y}}{\omega(\|x-y\|_{X})},$
and then define these three vanishing scales as
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{small}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to 0}\operatorname{osc}^{\omega}_{\delta}(f)=0\big{\}},\,$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{large}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\operatorname{osc}^{\omega}_{\delta}(f)=0\big{\}},$
$\dot{\operatorname{V}}^{0,\omega}_{\operatorname{far}}(X,Y)$ =$=\big{\{}f:X\to Y\,:\,\lim_{\delta\to\infty}\sup_{\min(\|x\|,\|y\|)>\delta}\operatorname{osc}^{\omega}_{(x,y)}(f)=0\big{\}}.$
Definition 1.1.
For each scale $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ set
$\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ $=\dot{\operatorname{V}}^{0,\omega}_{\Gamma}(X,Y)\cap\dot{C}^{0,\omega}(X,Y),$
$\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ $=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(X,Y).$
If $Y$ is a Banach space, then so are all of $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ as spaces of modulo additive constant equivalence classes; for full details, see Remark 2.1 below. From now on, whenever no confusion is possible, we start dropping the source and target spaces from the norms and simplify notation with $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|;$ in the same way, we write $\dot{C}^{0,\omega}(X,Y)=\dot{C}^{0,\omega}(X)=\dot{C}^{0,\omega},$ and so on.
We assume our moduli $\omega$ to be non-decreasing and satisfy
$\lim_{t\to 0}\omega(t)=0,\qquad\lim_{t\to\infty}\omega(t)=\infty,$ (2)
$\lim_{t\to 0}t/\omega(t)=0,$ (3)
$\omega(2t)\leq C_{\operatorname{db}}\omega(t),\qquad\mbox{for}\qquad t\in(0,\infty),$ (4)
where the doubling constant $C_{\operatorname{db}}>1.$ This is a big class of moduli that in particular covers the Hölder modulus $\alpha(t):=t^{\alpha},$ when $\alpha\in(0,1),$ but of course other more exotic moduli are also covered. The classes $\dot{\operatorname{VC}}^{0,\alpha}_{\operatorname{small}}$ are also called the little Hölder spaces. These are important in the study of Lipschitz algebras in metric spaces, and we refer to the monograph by Weaver [Weaverbook, Chapters 4 and 8] for a background. See also the recent monograph by D. Mitrea, I. Mitrea, and M. Mitrea [Mitreabook, Chapter 3] for a detailed exposition of the Hölder spaces in the setting of Ahlfors regular sets and their connection with functions of vanishing mean oscillation.
Notation
We denote $A\lesssim B$, if $A\leq CB$ for some constant $C>0$ depending only on an underlying function space that is fixed relative to ongoing considerations. For example, $C$ could depend on the modulus $\omega,$ or the dimension $n$ in the case $X=\mathbb{R}^{n}.$ Then $A\sim B$, if $A\lesssim B$ and $B\lesssim A.$ Subscripts or variables on constants and quantifiers $C_{a},C(a),\lesssim_{a}$ indicate their dependence on those subscripts.
1.2. Main results
We begin by stating a result without any smoothness on approximations by functions with bounded support. Many of our later results build on this one.
Theorem 1.2.
Let $X,Y$ be normed spaces, the modulus $\omega$ satisfy (2), (3) and (4). Then, there holds that
$\dot{\operatorname{VC}}^{0,\omega}(X,Y)=\overline{\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\cap C_{\operatorname{bs}}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}.$
Here and below, for any $\mathcal{C}(X,Y)\subset\dot{C}^{0,\omega}(X,Y),$ we denote by $\overline{\mathcal{C}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}$ the closure of $\mathcal{C}(X,Y)$ under the $\dot{C}^{0,\omega}(X,Y)$ seminorm. We refer here to the family of those $f\in\dot{C}^{0,\omega}(X,Y)$ for which there exists a sequence $\{f_{n}\}_{n}\subset\mathcal{C}(X,Y)$ so that $\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}(X,Y)}=0.$ The subscripted set $\mathcal{C}_{\operatorname{bs}}(X,Y)$ consists of boundedly supported functions in $\mathcal{C}(X,Y).$ In Theorem 1.2 $C_{\operatorname{bs}}(X,Y)$ denotes continuous functions with bounded support. In particular, the functions in $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ are continuous, and the intersection with $C_{\operatorname{bs}}(X,Y)$ gives the approximations bounded support.
While Theorem 1.2 does not provide any smoothness, it turns out to be a crucial step in our other theorems. For instance, letting $X=\mathbb{R}^{n}$ and $Y$ be a Banach space, Theorem 1.2 coupled with a mollification argument gives $C^{\infty}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$ approximations, this is Theorem 1.6 below. Both of these theorems will be proved in Section 2.
Our results are not confined to functions with a finite-dimensional source $X,$ and letting $Y=\mathbb{R},$ we prove smooth and Lipschitz approximation for the classes $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,\mathbb{R})$ and $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R}).$ (The corresponding approximations for $\mathbb{C}^{n}$-valued functions in place of $\mathbb{R}$-valued then follow after splitting into the coordinate functions and then each of these to the real and imaginary parts.) In particular, with $X$ being a separable normed space, we have the following general result.
Theorem 1.3.
Let $k\in\mathbb{N}\cup\{\infty\}$ and $X$ be a separable normed space admitting a $C^{k}$ and Lipschitz bump function. Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
  • (i)
    $\overline{C^{k}(X)\cap\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X),$
  • (ii)
    $\overline{C^{k}_{\operatorname{bs}}(X)\cap\mathrm{Lip}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}(X).$
Notice that, together with the $C^{k}$ order of smoothness, Theorem 1.3 guarantees Lipschitz approximations. In particular, any separable normed space $X$ with an equivalent norm of class $C^{k}$ satisfies the assumption of Theorem 1.3, see Remark 3.6. Moreover, when $X$ is a separable Hilbert space, one can arrange $C^{\infty}$ smooth approximations, thus obtaining an infinite dimensional version of Theorem 1.6, see Corollary 3.7. We will restate Theorem 1.3 and prove it in Subsection 3.1.
Another class of spaces $X$ on which we obtain smooth approximations are super-reflexive Banach spaces. A theorem of Pisier [Pisier] says that any super-reflexive space admits a renorming with modulus of smoothness of power type $1+\alpha,$ for some $\alpha\in(0,1]$ – see (35) below for the precise formulation. These spaces contain the Hilbert (separable or not) spaces, together with many of the classical Banach function spaces, such as the $L^{p}$ spaces, with $1<p<\infty;$ see Remark 3.11. In the following theorem we establish $C^{1,\alpha}$ and Lipschitz approximations for the classes $\dot{\operatorname{VC}}^{0,\omega}(X)$ and $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X).$
Theorem 1.4.
Let $X$ be a super-reflexive space that admits a renorming with modulus of smoothness of power type $1+\alpha,$ $\alpha\in(0,1].$ Then, for a modulus $\omega$ satisfying (2), (3) and (4), the following hold:
  • (i)
    $\overline{C^{1,\alpha}(X)\cap\mathrm{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$
  • (ii)
    $\overline{C^{1,\alpha}_{\operatorname{bs}}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}(X).$
In Remark 3.11, we will see that in Hilbert spaces and in $L^{p}$ spaces, with $p\geq 2,$ Theorem 1.4 yields $C^{1,1}$ approximations, i.e., $C^{1}$ functions with Lipschitz first derivatives.
Finally, for mappings $f:X\to Y,$ where $Y$ is an arbitrary Banach space, in Subsection 3.3 we obtain $C^{\infty}$ smooth approximations in the case where $X$ is a space of the form $X=c_{0}(\mathcal{A}).$
Theorem 1.5.
For an arbitrary set of indices $\mathcal{A},$ let $X=c_{0}(\mathcal{A}),$ and let $Y$ be a Banach space. Then, for any modulus $\omega$ satisfying (2), (3) and (4), the following hold:
  • (i)
    $\overline{C^{\infty}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y),$
  • (ii)
    $\overline{C^{\infty}_{\operatorname{bs}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}=\dot{\operatorname{VC}}^{0,\omega}(X,Y).$
Theorem 1.5 will be restated and proved in Subsection 3.3, see Theorem 3.12. There, we will also show that the approximations can be taken to be Lipschitz in the case $Y=\mathbb{R}.$
1.3. Applications on the Euclidean space
In the previous section we presented results that guarantee nice approximations, provided we know pointwise vanishing type conditions. When $X=\mathbb{R}^{n},$ we present formulations of these results in terms of mean oscillations type conditions, that are very convenient when studying the boundedness and compactness of commutators of Calderón-Zygmund operators (CZOs). In the one parameter setting our results recover important components to the theory, but from a more general approach. In the bi-parameter setting our results have very recently found applications by Martikainen and the second named author [MarOik24] in the study of the compactness of the bi-commutator. We describe all of this in full detail below, but to not get too far ahead of ourselves, let us begin by recapping boundedness and compactness characterizations of commutators in the one-parameter setting.
For a non-degenerate CZO $T$ on $\mathbb{R}^{n},$ a function $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C})$ (symbol of the commutator) and exponents $p,q\in(1,\infty),$ the commutator has the following mapping properties
$\|[b,T]\|_{L^{p}(\mathbb{R}^{n})\to L^{q}(\mathbb{R}^{n})}\sim\|b\|_{X^{p,q}(\mathbb{R}^{n})},$ (5)
where the space $X^{p,q}(\mathbb{R}^{n})$ is determined as follows. Given $p,q\in(1,\infty),$ denote by $\alpha(p,q)$ and $r(p,q)$ the exponents uniquely determined through the relations $\alpha(p,q)/n=1/p-1/q$ and $1/q=1/r(p,q)+1/p.$ Then, $X^{p,q}(\mathbb{R}^{n})$ is the space of bounded mean oscillations $\mathrm{BMO}(\mathbb{R}^{n}),$ when $p=q;$ the fractional BMO space $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n}),$ when $p<q;$ and the Lebesgue space $L^{r(p,q)}(\mathbb{R}^{n})$ modulo additive constants $\dot{L}^{r(p,q)}(\mathbb{R}^{n}),$ when $p>q.$ (The reader can recover the definitions of the spaces $\mathrm{BMO},\mathrm{BMO}^{\alpha}$ from Definition 1.7 below with the ``$\mbox{moduli}"$ $t\mapsto t^{0}$ and $t\mapsto t^{\alpha},$ respectively.) Emphasizing the recognition of the correct function spaces $X^{p,q}(\mathbb{R}^{n})$ to which the symbol belongs, the result (5) is due to Nehari [Nehari1957] ($n=1$) and Coifman, Rochberg, Weiss [CRW] ($n\geq 2$), when $q=p;$ due to Janson [Jan1978], when $q>p;$ and due to Hytönen [HyLpLq], when $q<p.$ We also refer the reader to the introductory section of [HyLpLq] for precise definitions and a fuller history of what was outlined above. Next, we describe the corresponding off-diagonal compactness characterization.
When $\alpha(p,q)>1,$ $\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n})$ consists of the constant functions and the commutator is bounded iff it is the zero operator (hence compact, in particular). Thus: for compactness only those exponents $p,q$ that result in $\alpha(p,q)\leq 1$ are interesting to study. When $\alpha(p,q)=1$ it is the case [GHWY21, Theorem 1.7. & 1.8.] that the commutator is compact iff it is the zero operator, again. Now for all $p,q\in(1,\infty)$ such that $\alpha(p,q)<1,$ there holds that
$[b,T]\in\mathcal{K}\left(L^{p}(\mathbb{R}^{n}),L^{q}(\mathbb{R}^{n})\right)\Leftrightarrow b\in Y^{p,q}(\mathbb{R}^{n}):=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{X^{p,q}(\mathbb{R}^{n})}.$ (6)
The characterization (6) is due to Uchiyama [Uch1978], when $p=q$; due to Guo, He, Wu and Yang [GHWY21], when $q>p;$ and due to Hytönen, Li, Tao and Yang [HLTY2023], when $q<p.$ When $q>p$ and $\alpha(p,q)\in(0,1),$ a result of Meyer’s [Mey1964] states that $\mathrm{BMO}^{\alpha(p,q)}(\mathbb{R}^{n})=\dot{C}^{0,\alpha(p,q)}(\mathbb{R}^{n}).$ Thus, the following Theorem 1.6 provides, in particular, a new characterization of the space $Y^{p,q}(\mathbb{R}^{n})$ in (6).
Theorem 1.6.
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2), (3) and (4). Then,
$\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)=\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$
To obtain the implications ``"``"`$\Rightarrow$"` ` ⇒ " in (6) it is very useful to have a description of $Y^{p,q}(\mathbb{R}^{n}),$ in the cases $q\geq p,$ in terms of some vanishing mean oscillation (VMO) criteria. Such VMO criteria (Definition 1.8, below) usually follow immediately from the existence of approximations, but to establish approximations from VMO criteria is delicate. Recall that in the previous section we established existence of approximations beginning from some pointwise vanishing criteria $\dot{\operatorname{VC}}.$ Below, as Theorem 1.9, we provide the correspondence between these $\mathrm{VMO}$ and the $\dot{\operatorname{VC}}$ type criterions, which allows us, when $X=\mathbb{R}^{n},$ to reformulate all of our results, among them Theorem 1.6, in terms of mean oscillations.
1.4. Formulation as mean oscillation conditions
Definition 1.7.
Let $\omega$ be a modulus and $Y$ a Banach space. Then, $\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)$ consists of those $f\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},Y)$ locally Bochner integrable functions for which
$\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n})}=\sup_{Q}\mathcal{O}^{\omega}(f;Q)<\infty,\qquad\mathcal{O}^{\omega}(f;Q)=\frac{1}{\omega(\ell(Q))}\fint_{Q}|f-\langle f\rangle_{Q}|_{Y},$
where the supremum is taken over all cubes in $\mathbb{R}^{n}.$
Definition 1.8.
Let $\omega$ be a modulus and $Y$ a Banach space and define the vanishing scales
$\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to 0}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\ell(Q)=\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
$\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)$ =$=\Big{\{}f\in\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)\,:\,\lim_{\delta\to\infty}\sup_{\begin{subarray}{c}\operatorname{\mathit{d}}(Q,0)>\delta\end{subarray}}\mathcal{O}^{\omega}(f;Q)=0\Big{\}},$
where the suprema are taken over all cubes. Then, define
$\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y)$ $=\mathrm{VMO}^{\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y)\cap\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)\cap\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$
We continue by noting that the approximation $\mathrm{VMO}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}(\mathbb{R}^{n})}$ was already present in the classical work of Uchiyama [Uch1978]. Uchiyama’s construction does not easily translate to the case $\alpha\in(0,1)$ and thus in [GHWY21] a non-trivial approximation for $\mathrm{VMO}^{\alpha}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\mathrm{BMO}^{\alpha}(\mathbb{R}^{n})}$ was provided that relies, more or less, on the existence of dyadic grids on the underlying space $\mathbb{R}^{n}$ (as does Uchiyama’s original approximation) and a careful construction using these grids. Such a construction, when $X$ is an arbitrary normed space, has no hope of working, simply because there is no grid available. Therefore, one more fact of our Theorem 1.6, in particular of its proof, is that it provides a conceptually easier approximation that uses only basic mollification and a carefully chosen truncation of the support. To make the connection with the mean oscillations and the pointwise conditions, we assume the following summability condition
$[\omega]_{*}:=\sup_{s>0}\frac{1}{\omega(s)}\int_{0}^{s}\omega(t)\frac{\,\mathrm{d}t}{t}<\infty.$ (7)
Theorem 1.9.
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (4) and (7). Then,
$\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)},\qquad\|\cdot\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}\lesssim_{n,C_{\operatorname{db}}}[\omega]_{*}\|\cdot\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$ (8)
Moreover, if in addition $\omega(\infty)=\infty$, then for each scale $\Gamma\in\{\operatorname{small},\operatorname{far},\operatorname{large}\},$ we have
$\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y),\qquad\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y).$ (9)
The proof-idea of Theorem 1.9 goes back to Meyers [Mey1964], who obtained a version of the bounds (8) with the modulus $t^{\alpha},$ for $\alpha>0.$ The condition (7) appears new, but we recognize that it is a natural summability condition, after all. The identification (9) is completely new. Combining Theorems 1.6 and 1.9, we immediately obtain the following.
Theorem 1.10.
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2), (3), (4) and (7). Then,
$\mathrm{VMO}^{\omega}(\mathbb{R}^{n},Y)=\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
1.5. Applications to compactness of the bi-commutator
Banach space valued operators and function spaces often appear in bi-parameter harmonic analysis. Then, it is maybe no surprise that the Banach valued Theorems (1.6), 1.10 and 1.9 have already found applications to bi-parameter harmonic analysis in the recent work [MarOik24]. We turn to the details.
We now consider $\mathbb{R}^{n}=\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$ as a bi-parameter space, take non-degenerate CZOs $T_{i}$ on $\mathbb{R}^{n_{i}}$ and $b\in L^{1}_{\operatorname*{loc}}(\mathbb{R}^{n},\mathbb{C}).$ Then, the operator of interest, the bi-commutator, is
$[T_{2},[b,T_{1}]]=T_{2}[b,T_{1}]-[b,T_{1}]T_{2}=T_{2}bT_{1}-T_{2}T_{1}b-bT_{1}T_{2}+T_{1}bT_{2},$
where $T_{i}$ acts on the variable $x_{i}.$ The product BMO upper bound for the $L^{p}\to L^{p}$ boundedness of the bi-commutator was first obtained with the Hilbert transforms in Ferguson, Sadosky [FerSad], and in full generality of completely arbitrary CZOs later by Dalenc, Ou [DaO16]. The corresponding lower bound in terms of product $\mathrm{BMO}$ for special singular integrals (Hilbert/Riesz transforms) has a somewhat complicated history. While the product $\mathrm{BMO}$ is still believed to be the characterizing condition, the existing proofs have been reported to have a gap, see e.g. [Ferguson2002, HTV2021ce, AHLMO2021], and to the best of our knowledge a fix has not yet been published. The existence of this gap adds considerable interest to all natural questions involving the bi-commutator.
Now, in place of considering $L^{p}$-to-$L^{p}$ bounds it is natural to consider mixed $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ bounds, where we allow $q_{i}\not=p_{i},$ this was done recently in [AHLMO2021, Theorem 1.1.]. The exponents that directly connect with our work are those for which $\beta_{i}:=n_{i}(1/p_{i}-1/q_{i})\geq 0$ and $\beta_{i}>0$ for some $i=1,2,$ which we assume here and below. Then, the bi-parameter space to characterize mixed $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ boundedness turned out to be
$\|b$ $\|_{\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})}=\sup_{R=I_{1}\times I_{2}}\mathcal{O}^{\beta_{1},\beta_{2}}(b,I_{1}\times I_{2})$
$=\sup_{R=I_{1}\times I_{2}}\ell(I_{1})^{-\beta_{1}}\fint_{I_{1}}\ell(I_{2})^{-\beta_{2}}\fint_{I_{2}}|b(x_{1},x_{2})-\langle b(x_{1},\cdot)\rangle_{I_{2}}-\langle b(\cdot,x_{2})\rangle_{I_{1}}+\langle b\rangle_{I_{1}\times I_{2}}|\,\mathrm{d}x_{2}\,\mathrm{d}x_{1}.$
The space $\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})$ is in fact a $\mathrm{BMO}$-valued or $\dot{C}^{0,\alpha}$-valued Hölder space , see [AHLMO2021, Section 3]. Then, very recently [MarOik24, Theorem 1.1.] the corresponding $L^{p_{1}}(L^{p_{2}})$-to-$L^{q_{1}}(L^{q_{2}})$ compactness was characterized through the VMO analogue $\operatorname{VMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n}).$ Moreover, as a corollary of these truly mixed norm compactness characterizations the natural product BMO condition in the non-mixed $L^{p}\to L^{p}$ case was attained as a sufficient condition. One of the crucial theorems in [MarOik24] is the following theorem, which is, in fact, a corollary of the approximations developed in this article.
Theorem 1.11.
Let $\beta_{i}\in[0,1)$ and $\beta_{i}>0$ for some $i=1,2,$ then
$\operatorname{VMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})=\overline{C^{\infty}_{c}(\mathbb{R}^{n})}^{\operatorname{BMO}^{\beta_{1},\beta_{2}}(\mathbb{R}^{n})}.$
Theorem 1.11 is a bi-parameter analogue of Theorem 1.10 (with $Y=\mathbb{C}$), where in at least one parameter we have a vanishing Hölder type behaviour. The proof of Theorem 1.11 is split between the current article and [MarOik24]. Indeed, we record here a quantitative version of Theorem 1.6 in the body-text after the proof of Theorem 1.6 as Corollary 2.5; then, assuming Corollary 2.5, Theorem 1.11 is proved in [MarOik24, Section 3.].
Acknowledgements
C. M. was supported by the grant no. 334466 of the Research Council of Norway, “Fourier Methods and Multiplicative Analysis”. T.O. was supported by the Finnish Academy of Science and Letters, and by the MICINN (Spain) grant no. PID2020-114167GB-I00.
2. Bounded support approximation
In this section we prove Theorem 1.2 and then show how to obtain from this Theorem 1.6 and then Corollary 2.5. After these, at the end of this section, we prove Theorem 1.9. Recall that we simplify notation by denoting $\|x\|_{X}=\|x\|$ and $\|f(x)\|_{Y}=|f(x)|.$ We begin with a simple remark.
Remark 2.1.
For normed spaces $X$ and $Y,$ and $\Gamma\in\{\operatorname{small},\operatorname{large},\operatorname{far}\},$ the set $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ is closed with respect to the $\dot{C}^{0,\omega}$-seminorm, meaning that if a sequence of functions $(f_{n})_{n}\subset\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ converges to an $f\in\dot{C}^{0,\omega}(X,Y),$ then $f\in\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y)$ as well. Indeed, for any $\varepsilon>0,$ we find $n\in\mathbb{N}$ so that $\|f-f_{n}\|_{\dot{C}^{0,\omega}}\leq\varepsilon$. Then, for any two distinct points $x,y\in X,$ we write
$\frac{|f(x)-f(y)|}{\omega(\|x-y\|)}$ $\leq\frac{|f_{n}(x)-f_{n}(y)|}{\omega(\|x-y\|)}+\frac{|(f-f_{n})(x)-(f-f_{n})(y)|}{\omega(\|x-y\|)}$
$\leq\frac{|f_{n}(x)-f_{n}(y)|}{\omega(\|x-y\|)}+\|f-f_{n}\|_{\dot{C}^{0,\omega}(X,Y)}\leq\frac{|f_{n}(x)-f_{n}(y)|}{\omega(\|x-y\|)}+\varepsilon.$
Since $f_{n}\in\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y),$ the above clearly implies $f\in\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y).$
Consequently, in the case where $Y$ is a Banach space, $(\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(X,Y),\|\cdot\|_{\dot{C}^{0,\omega}(X,Y)})$ is a Banach space of modulo constant equivalence classes.
The following two lemmas will be very useful.
Lemma 2.2.
Let $X,Y$ be arbitrary normed spaces. Let $\omega$ satisfy (2) and (4). Then, we have
$\dot{\operatorname{VC}}_{\operatorname{far}}^{0,\omega}(X,Y)=\Big{\{}f\in\dot{C}^{0,\omega}(X,Y)\,:\,\lim_{\delta\to\infty}\sup_{\max(\|x\|,\|y\|)>\delta}\operatorname{osc}^{\omega}_{(x,y)}(f)=0\Big{\}}.$
Proof.
Only the inclusion $\subset$ is not immediate; let $f\in\dot{\operatorname{VC}}_{\operatorname{far}}^{0,\omega}(X,Y)$ and $\varepsilon>0.$ Let $M=M(\varepsilon)$ be such that if $y,z\in X$ and $\|y\|,\|z\|\geq M,$ then $|f(y)-f(z)|\leq\varepsilon\omega(\|y-z\|).$ Now we consider arbitrary $x,y\in X,$ and assume $\|y\|>R$, for certain $R\gg M$ to be specified later. If $\|x\|>M,$ then we are done by how we fixed $M$ above. So suppose that $x\in B(0,M).$ Take a point $z\in[x,y]$ with $\|z\|=M.$ By $\|x\|,\|z\|\leq M$ and $\|y\|,\|z\|\geq M,$
$\begin{split}\frac{|f(x)-f(y)|}{\omega(\|x-y\|)}&\leq\frac{|f(x)-f(z)|}{\omega(\|x-y\|)}+\frac{|f(z)-f(y)|}{\omega(\|z-y\|)}\cdot\frac{\omega(\|z-y\|)}{\omega(\|x-y\|)}\\ &\leq\frac{\|f\|_{\dot{C}^{0,\omega}(X)}\omega(\|x-z\|)}{\omega(\|x-y\|)}+\varepsilon\leq\frac{\|f\|_{\dot{C}^{0,\omega}(X)}\omega(2M)}{\omega(\|x-y\|)}+\varepsilon.\end{split}$ (10)
To control the above term, note that if $\|y\|>R\gg M,$ then we have $\|x-y\|\geq R/2$ because $\|x\|\leq M.$ Also, by condition $\lim_{t\to\infty}\omega(t)=\infty$, we find $R=R(M,\varepsilon)$ so that $\omega(2M)\leq\varepsilon\omega(R)$. Therefore, since $\|y\|\geq R,$ using the doubling condition (4) we conclude
[…]
Lemma 2.3.
Let $X,Y$ be arbitrary normed spaces. Let $\omega$ be non-decreasing and satisfy (4). Let $\tau:X\to X$ be Lipschitz. Then,
$\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)\circ\tau\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y).$
Proof.
Let $\varepsilon>0$ and we need to show that if $r$ is taken sufficiently small, then
$\|f\circ\tau\|_{\dot{C}^{0,\omega}_{r}(X,Y)}:=\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|<r\end{subarray}}\frac{|(f\circ\tau)(x)-(f\circ\tau)(y)|}{\omega(\|x-y\|)}\leq\varepsilon.$ (11)
Let us denote $\varepsilon(r):=\|f\|_{\dot{C}^{0,\omega}_{r}}$ so that $\varepsilon(r)\to 0$ as $r\to 0.$ Let $\|x-y\|<r$ and by $\omega$ being non-decreasing and doubling we bound
$\begin{split}\frac{|(f\circ\tau)(x)-(f\circ\tau)(y)|}{\omega(\|x-y\|)}&=\frac{|f(\tau(x))-f(\tau(y))|}{\omega(\|\tau(x)-\tau(y)\|)}\frac{\omega(\|\tau(x)-\tau(y)\|)}{\omega(\|x-y\|)}\\ &\leq\varepsilon(\operatorname{Lip}(\tau)r)\cdot C(C_{\operatorname{db}},\operatorname{Lip}(\tau)),\end{split}$
and the right-hand side tends to zero as $r$ does. ∎
Proof of Theorem 1.2.
First, note that if $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and has bounded support, then, by property (3) of $\omega,$ $f$ belongs to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ Since $\dot{\operatorname{VC}}^{0,\omega}(X,Y)$ is closed under limits with respect to the seminorm $\dot{C}^{0,\omega},$ as per Remark 2.1, the inclusion ``"``"`$\subset$"` ` ⊂ " holds true.
To prove the reverse (much more complicated) inclusion, let $f\in\dot{\operatorname{VC}}^{0,\omega}$ and $\varepsilon>0.$ By Lemma 2.2 we find two parameters $0<r\ll R$ such that
$\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|\leq r\end{subarray}}\frac{\left|f(x)-f(y)\right|}{\omega(\|x-y\|)}\leq\varepsilon,$ (12)
and
$\sup_{x\in B(0,R)^{c}}\sup_{y\in X}\frac{\left|f(x)-f(y)\right|}{\omega(\|x-y\|)}\leq\varepsilon.$ (13)
We consider an auxiliary parameter $M\gg R$ and the following function
$\tau=\tau_{M}:X\to B(0,M),\qquad\tau(x)=\begin{cases}x,\qquad&\|x\|<M,\\ \left(\frac{2M-\|x\|}{M}\right)^{2}x,\qquad M\leq&\|x\|<2M,\\ 0,\qquad&\|x\|\geq 2M.\end{cases}$ (14)
We next show that if $r$ is small enough and $R,M$ are large enough then
$f\circ\tau\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}},\qquad\|f-f\circ\tau\|_{\dot{C}^{0,\omega}}\lesssim\varepsilon.$ (15)
Notice that $f\circ\tau-f(0)$ is zero outside $B(0,2M),$ and hence $f\circ\tau$ has bounded support. (Recall that $\dot{C}^{0,\omega}$ is defined modulo constant equivalence classes.) To actually verify the claims of (15) we begin checking some properties of the function $\tau.$
First, we verify
$\|\tau(x)-\tau(z)\|\leq 5\|x-z\|\quad\text{for all}\quad x,z\in X,$ (16)
and in particular by doing so, Lemma 2.3 implies the left claim on the line (15). Indeed, for points $x,z\in B(0,2M)\setminus B(0,M),$ the definition of $\tau$ and the triangle inequality give
$\begin{split}\|\tau(x)-\tau(z)\|&\leq\Big{|}\big{(}2-\frac{\|x\|}{M}\big{)}^{2}-\big{(}2-\frac{\|z\|}{M}\big{)}^{2}\Big{|}\|z\|+\big{(}2-\frac{\|x\|}{M}\big{)}^{2}\|x-z\|\\ &\leq\big{(}2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\big{)}\frac{\|x-z\|}{M}\|z\|+\big{(}2-\frac{\|x\|}{M}\big{)}^{2}\|x-z\|\\ &\leq\big{(}4-\frac{2M}{M}\big{)}\frac{\|x-z\|}{M}(2M)+\big{(}2-\frac{M}{M}\big{)}^{2}\|x-z\|=5\|x-z\|,\end{split}$ (17)
where in the second bound we used the basic identity $a^{2}-b^{2}=(a-b)(a+b).$ As $\tau$ is obviously $1$-Lipschitz in the sets $B(0,M)$ and $B(0,2M)^{c},$ and $\tau$ is continuous in $X,$ we deduce (16) for all $x,z\in X.$
Then we show the following contraction property:
$\begin{split}\Big{\{}x\in B(0,M)^{c},\quad\tau(x),\tau(z)\in B(0,R)\Big{\}}\Longrightarrow\|\tau(x)-\tau(z)\|\lesssim R\,\theta(M)\|x-z\|,\end{split}$ (18)
where $\theta:(0,\infty)\to(0,\infty)$ is a function satisfying $\lim_{M\to\infty}\theta(M)=0$, whose definition will be specified while we next check the truth of the implication [^1] We remark that (18) would not be true, if we replaced the power $2$ in the definition (14) of $\tau$ by the power $1$. If $\|x-z\|>M/2,$ then we have
$\|\tau(x)-\tau(z)\|\leq 2R<2R\frac{\|x-z\|}{M/2}=4R\,\theta_{1}(M)\|x-z\|,\qquad\theta_{1}(M):=\frac{1}{M}.$
So next assume that $z\in B(x,M/2).$ Since $R\ll M$ and $\tau(z)\in B(0,R)$, by the definition of $\tau$ we must have necessarily $\|z\|>M.$ Thus, it remains to show that
$\begin{split}\Big{\{}x,z\in B(0,M)^{c},\quad\tau(x),\tau(z)\in B(0,R)\Big{\}}\Longrightarrow\|\tau(x)-\tau(z)\|\lesssim R\,\theta(M)\|x-z\|,\end{split}$ (19)
which is symmetric with respect to both variables $x,z.$ Notice that if $x,z\in B(0,2M)^{c},$ then the left-hand side of the claimed estimate is zero. Assume that $x,z\in B(0,2M)\setminus B(0,M)$ and we make the following observation. By $\tau(x)\in B(0,R)$ and $x\in B(0,M)^{c},$ there holds that
$\Big{\|}\left(\frac{2M-\|x\|}{M}\right)^{2}x\Big{\|}\leq R\Longrightarrow\Big{|}\frac{2M-\|x\|}{M}\Big{|}\leq\sqrt{\frac{R}{\|x\|}}\leq\sqrt{\frac{R}{M}}=\sqrt{R}\,\theta_{2}(M),\qquad\theta_{2}(M):=\frac{1}{\sqrt{M}},$
and the same bound is valid for the variable $z.$ Thus, by (17) we obtain
$\|\tau(x)-\tau(z)\|$ $\leq\Big{(}2-\frac{\|x\|}{M}+2-\frac{\|z\|}{M}\Big{)}\frac{\|x-z\|}{M}\|z\|+\Big{(}2-\frac{\|x\|}{M}\Big{)}^{2}\|x-z\|$
$\leq\Big{(}4\sqrt{\frac{R}{M}}+\frac{R}{M}\Big{)}\|x-z\|\leq 5R\,\theta_{2}(M)\|x-z\|.$
As the last case, suppose that $z\in B(0,2M)^{c}.$ This case follows by the continuity of $\tau$ and the previous case. Indeed, taking a point $y\in[x,z]$ with $\|y\|=2M,$ by the previous case applied for $x$ and $y,$ and bearing in mind that $\tau(z)=\tau(y)=0,$ we deduce
$\|\tau(x)-\tau(z)\|=\|\tau(x)-\tau(y)\|\leq 5R\,\theta_{2}(M)\|x-y\|\leq 5R\,\theta_{2}(M)\|x-z\|.$
All in all, we have now shown that (18) is valid with the function $\theta:(0,\infty)\to(0,\infty)$ given by
$\theta(M):=\frac{1}{\sqrt{M}}\geq\frac{1}{2}(\theta_{1}(M)+\theta_{2}(M)).$ (20)
Now, let us denote $g:=f\circ\tau\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ and let $r^{\prime}>0$ be such that
$\|g\|_{\dot{C}^{0,\omega}_{r^{\prime}}}:=\sup_{\begin{subarray}{c}x\not=y\in X\\ \|x-y\|<r^{\prime}\end{subarray}}\frac{|g(x)-g(y)|}{\omega(\|x-y\|)}\leq\varepsilon.$ (21)
By (12), the fact that $\operatorname{Lip}(\tau)\leq 5$ (see (16)), and the proof of Lemma 2.3, this new $r^{\prime}>0$ can be taken to be an absolute multiple of the $r$ in (12). Therefore, we may and do assume that both (12) and (21) hold for the same $r=r(\varepsilon),$ which is of course independent of $R$ and $M.$
Next we show that
$\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|g(x)-g(z)\right|}{\omega(\|x-z\|)}\lesssim\varepsilon,$ (22)
provided $M$ is large enough. By (21), it is enough to estimate (22) only at those couples $x,z$ with $\|x-z\|\geq r.$ Fix a point $x\in B(0,M)^{c}$ and we distinguish into cases. If $\tau(x)\in B(0,R)^{c},$ then by (13) and $\omega$ being non-decreasing and doubling we obtain
$\begin{split}\left|g(x)-g(z)\right|&=\left|f(\tau(x))-f(\tau(z))\right|\\ &\leq\Big{(}\sup_{u\in B(0,R)^{c}}\sup_{v\in X}\frac{\left|f(u)-f(v)\right|}{\omega(\|u-v\|)}\Big{)}\omega(\|\tau(x)-\tau(z)\|)\lesssim_{C_{\operatorname{db}}}\varepsilon\omega(\|x-z\|).\end{split}$ (23)
The case $\tau(z)\in B(0,R)^{c}$ is symmetrical.
Now consider $\tau(x),\tau(z)\in B(0,R)$ and we split into subcases. But first observe that for those $x,z$ with $\|x-z\|\geq M^{1/4},$ using that $f\in\dot{C}^{0,\omega}$ we can write
$\left|g(x)-g(z)\right|=\left|f(\tau(x))-f(\tau(z))\right|\leq\|f\|_{\dot{C}^{0,\omega}}\omega(|\tau(x)-\tau(z)|)\leq\|f\|_{\dot{C}^{0,\omega}}\omega(2R).$
Provided that $M=M(\varepsilon,R)$ is large enough, the condition $\omega(\infty)=\infty$ says that the last term can be made smaller than $\varepsilon\omega(M^{1/4})\leq\varepsilon\omega(\|x-z\|),$ and thus (22) holds for these $x$ and $z.$ Therefore, during the rest of the verification of (22) we can assume that $\|x-z\|\leq M^{1/4},$ whenever needed.
Notice also that in the case $x,z\in X\setminus B(0,2M)$ there is nothing to prove, as then $g(x)=g(z)=0.$
Thus, we can assume that both $x,z\in B(0,10M).$ In this case we bound
$\begin{split}\left|g(x)-g(z)\right|&=\left|f(\tau(x))-f(\tau(z))\right|\\ &\lesssim\|f\|_{\dot{C}^{0,\omega}(B(0,2R))}\omega(\|\tau(x)-\tau(z)\|)\lesssim\omega(\|\tau(x)-\tau(z)\|).\end{split}$ (24)
Let $\theta_{x,z}(M)$ be defined by the identity
$\|\tau(x)-\tau(z)\|=:\theta_{x,z}(M)\|x-z\|,\qquad\theta_{x,z}(M)\lesssim R/\sqrt{M},$
where the bound follows by (18) and (20). By the condition $\lim_{t\to 0}\omega(t)=0$, we choose $M=M(\varepsilon,r,R)$ sufficiently large so that $\omega(RM^{-1/4})\leq\varepsilon\omega(r)$. Since we are assuming $r\leq\|x-z\|\leq M^{1/4},$ we obtain
[…]
Chaining the estimates (24) and (25), we have shown (22) for $x,z$ as in the present case.
Finally we consider the case $\tau(x),\tau(z)\in B(0,R),$ and $x\in B(0,2M)\setminus B(0,M)$ and $z\in X\setminus B(0,10M).$ We pick an auxiliary point $y$ such that $|y|=2M.$ Using $\tau(z)=\tau(y)=0,$ the bound (25) (valid by $x,y\in B(0,10M)$) and that $\omega$ is non-decreasing we obtain
$\begin{split}|g(x)-g(z)|=|g(x)-g(y)|\leq\|f\|_{\dot{C}^{0,\omega}(X)}\omega(\|\tau(x)-\tau(y)\|)\lesssim\varepsilon\omega(\|x-y\|)\leq\varepsilon\omega(\|x-z\|),\end{split}$
which is a bound of the correct form. The case $\tau(x),\tau(z)\in B(0,R)$ and $z\in B(0,2M)$ and $x\in X\setminus B(0,10M)$ is symmetrical to the last one.
We now fix the parameters $M\gg R\gg r$ so that both (21) and (22) hold. Then, we show the right claim on the line (15). If $x,z\in B(0,M)$ there is nothing to prove, since $\tau(x)=x$ and $\tau(z)=z.$ Finally, we check that
$\sup_{x\in B(0,M)^{c}}\sup_{z\in\mathbb{R}^{n}}\frac{\left|(f-g)(x)-(f-g)(z)\right|}{\omega(\|x-z\|)}$
$\qquad\leq\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|f(x)-f(z)\right|}{\omega(\|x-z\|)}+\sup_{x\in B(0,M)^{c}}\sup_{z\in X}\frac{\left|g(x)-g(z)\right|}{\omega(\|x-z\|)}\lesssim\varepsilon,$
and hence have now shown both claims on the line (15). ∎
We next show how to easily upgrade the above bounded support approximation into a compact and smooth approximation on $X=\mathbb{R}^{n},$ but $Y$ is allowed to be an arbitrary Banach space.
Proof of Theorem 1.6.
Let $f\in\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ and by Theorem 1.2 find $g\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(\mathbb{R}^{n},Y)$ with bounded support such that $\|f-g\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}<\varepsilon.$ Let $r>0$ and $0\leq\eta_{r}\in C^{\infty}_{c}(B(0,r))$ be a standard real-valued smooth bump in $\mathbb{R}^{n}$ with $\int_{\mathbb{R}^{n}}\eta_{r}=1,$ and we define $h_{r}:=g*\eta_{r},$ where the integral is understood in the Bochner sense. Notice that $h_{r}$ is compactly supported, as $g$ is, and smooth $h_{r}\in C^{\infty}(\mathbb{R}^{n},Y)$ as a smooth mollification. Provided that $|u|\leq\delta,$ there holds uniformly in $r$ that
$|h_{r}(x+u)-h_{r}(x)|=\Big{|}\int_{\mathbb{R}^{n}}\eta_{r}(y)(g(x+u+y)+g(x+y))\,\mathrm{d}y\Big{|}\leq\|g\|_{\dot{C}^{0,\omega}_{\delta}}\omega(|u|),$
and hence we find $\delta>0$ so that, uniformly in $r,$ we have $\|g-h_{r}\|_{\dot{C}^{0,\omega}_{\delta}}\leq\|g\|_{\dot{C}^{0,\omega}_{\delta}}+\|h_{r}\|_{\dot{C}^{0,\omega}_{\delta}}\leq\varepsilon.$ To deal with the scales $\geq\delta$ we argue as follows. Observe that $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ is uniformly continuous, by $\lim_{t\to 0}\omega(t)=0,$ and thus the approximation $h_{r}:=g*\eta_{r}$ converges uniformly to $g,$ as $r\to 0.$ We let $r=r(\delta)$ be so small that $\|g-h_{r}\|_{\infty}\leq\frac{1}{2}\varepsilon\omega(\delta).$ Then, for $|x-y|>\delta,$ there holds that
$\|(g-h_{r})(x)-(g-h_{r})(y)\|\leq 2\|g-h_{r}\|_{\infty}\leq\varepsilon\omega(\delta)\leq\varepsilon\omega(|x-y|),$
by $\omega$ being non-decreasing. This concludes the proof of $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)\subset\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}.$
For the reverse inclusion ``","```$\supset$",` ` ⊃ " , simply notice that by the Mean Value Inequality, $C^{\infty}_{c}(\mathbb{R}^{n},Y)\subset\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$. We remind that this inequality holds for smooth $Y$-valued functions, as a consequence of the Hahn-Banach theorem on $Y.$ Also, by (3) there holds that $\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)\subset\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$ As $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ is closed with respect to $\dot{C}^{0,\omega}$ limits, by Remark 2.1, we are done. ∎
Next, we record a quantitative version of Theorem 1.6 to be used in the proof of Theorem 1.11 in [MarOik24]. While Corollary 2.5 below is one the key elements in applications to the bi-commutator in [MarOik24], it plays no further role in the present article.
Definition 2.4.
For a subset $G$ of $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$ we write $G\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ (with ``$u"$ for uniformly) provided that $G$ is equibounded in the $\dot{C}^{0,\omega}$ seminorm, and that for all $\varepsilon>0,$ there exists $t>0$ such that if $|x-y|<t$ or $|x|>t^{-1}$ or $|y|>t^{-1},$ then
$\sup_{g\in G}\frac{|g(x)-g(y)|}{\omega(|x-y|)}<\varepsilon.$
By an inspection of the above proofs, the reader convinces themselves that the following is true.
Corollary 2.5.
Let $Y$ be a Banach space and the modulus $\omega$ satisfy (2) and (4). Let $G\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$ Then,
  • for all $L>0$ we have $G\circ\tau_{L}\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$
  • for all $L,L^{\prime}>0$ we have $(G\circ\tau_{L})*\eta_{L^{\prime}}\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$
  • for all $M=M(\varepsilon)>0$ sufficiently large
    $\sup_{f\in G}\|f-f\circ\tau_{M}\|_{\dot{\operatorname{C}}^{0,\omega}(\mathbb{R}^{n},Y)}\leq\varepsilon,$ (26)
  • for all $K=K(M,\varepsilon)>0$ sufficiently large
    $\sup_{f\in G}\|f\circ\tau_{M}-(f\circ\tau_{M})*\eta_{1/K}\|_{\dot{\operatorname{C}}^{0,\omega}(\mathbb{R}^{n},Y)}\leq\varepsilon.$ (27)
Proving approximations with a more general normed space $X$ and $Y=\mathbb{R}$ (or $Y=\mathbb{C}$) is a delicate task and this is the content of the next Section 3. But before that, we finish this section by proving the connection between pointwise and mean oscillation type conditions.
Proof of Theorem 1.9.
We first prove that the norms are equivalent. So fix an arbitrary cube $Q\in\mathcal{Q}$ and estimate
$\fint_{Q}|f(x)-\langle f\rangle_{Q}|_{Y}\,\mathrm{d}x\leq\fint_{Q}\fint_{Q}|f(x)-f(y)|_{Y}\,\mathrm{d}x\,\mathrm{d}y\leq\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\fint_{Q}\fint_{Q}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y$
and we continue the bound with
$\fint_{Q}\fint_{Q}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y\lesssim_{n,C_{\operatorname{db}}}\omega(\ell(Q)),$ (28)
thus we obtain $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\gtrsim_{n,C_{\operatorname{db}}}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
For the other direction, let $x,z\in\mathbb{R}^{n}$ be arbitrary and let $Q_{0}\subset\mathbb{R}^{n}$ be a cube containing both $x$ and $y$ such that $\ell(Q_{0})=|x-y|.$ Consider the dyadic descendants of $Q_{0}$ (achieved by iteratively halving the sides) and for every $k\in\mathbb{N}\cup\{0\}$ let $Q_{k}(x)$ be the descendant of $Q_{0}$ of sidelength $2^{-k}\ell(Q_{0})$ that contains the point $x.$ Similarly, we define $Q_{k}(y)$ for each $k.$ By the continuity of $f:\mathbb{R}^{n}\to Y,$ we write
$f(x)-f(y)=\Big{(}\sum_{k=0}^{\infty}\langle f\rangle_{Q_{k+1}(x)}-\langle f\rangle_{Q_{k}(x)}\Big{)}-\Big{(}\sum_{k=0}^{\infty}\langle f\rangle_{Q_{k+1}(y)}-\langle f\rangle_{Q_{k}(y)}\Big{)}.$ (29)
Both sums are estimated identically, so consider the first:
$\Big{|}\sum_{k=0}^{\infty}\langle f\rangle_{Q_{k+1}(x)}-\langle f\rangle_{Q_{k}(x)}\Big{|}_{Y}$ $\lesssim_{n}\sum_{k=0}^{\infty}\fint_{Q_{k}(x)}|f-\langle f\rangle_{Q_{k}(x)}|_{Y}\leq\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}\sum_{k=0}^{\infty}\omega(\ell(Q_{k}(x)))$
and we continue the bound, using that $\omega$ is non-decreasing in the first bound, with
$\sum_{k=0}^{\infty}\omega(\ell(Q_{k}(x)))=\sum_{k=0}^{\infty}\ell(Q_{k}(x))\frac{\omega(\ell(Q_{k}(x)))}{\ell(Q_{k}(x))}\leq 2\int_{0}^{\ell(Q_{0})}\frac{\omega(t)}{t}\,\mathrm{d}t\lesssim[\omega]_{*}\omega(\ell(Q_{0})),$
and, since $\omega(\ell(Q_{0}))=\omega(|x-y|),$ we have shown that $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n}[\omega]_{*}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
It is clear that the above argument gives $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ when $\Gamma=\operatorname{small},$ as well as the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ for $\Gamma=\operatorname{far}.$ The reverse inclusion ``""```$\supset$"` ` ⊃ " for $\Gamma=\operatorname{far}$ can be seen as follows. Let $R>0$ be such that if $\operatorname{\mathit{d}}(Q,0)>R,$ then $\mathcal{O}^{\omega}(f,Q)\leq\varepsilon.$ Let $|x|,|y|>R$ both be far from the origin. Using the structure of the Euclidean space, clearly there exists a dimensional constant $N(n)$ and points and cubes $\{z_{k},Q_{k}\}_{k=1}^{N}$ such that $z_{0}=x,$ $z_{N}=y$ and $z_{k},z_{k+1}\in Q_{k}$ and $\operatorname{\mathit{d}}(Q_{k},0)\geq R$ and $\ell(Q_{k})=|z_{k}-z_{k+1}|.$ Thus, by the chain of inequalities that led us to the right estimate for (29), we obtain
$|f(x)-f(y)|_{Y}\leq\sum_{k=1}^{N}|f(z_{k})-f(z_{k+1})|_{Y}\lesssim[\omega]_{*}\varepsilon\sum_{k=1}^{N}\omega(\ell(Q_{k}))\lesssim\varepsilon\omega(|x-y|),$
which concludes the proof of the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)\supset\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y).$
For $\Gamma=\operatorname{large},$ a further argument is needed for both inclusions. We begin by showing that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\supset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y),$ and $\varepsilon>0,$ and let $K\in\mathbb{N}$ be such that if $\ell(Q)\geq 2^{K},$ then $\mathcal{O}^{\omega}(f;Q)\leq\varepsilon.$ Let $N\gg K$ be a large integer to be specified later, and suppose $x,y\in\mathbb{R}^{n}$ are so that $|x-y|\geq 2^{N}.$ Let $M\geq N$ be so that $2^{M}\leq\left|x-y\right|\leq 2^{M+1},$ and let $Q_{0}$ be a cube containing both $x,y$ and with $\ell(Q_{0})=|x-y|.$ Then, again considering the left sum in the expansion (29), we bound
$\Big{|}\sum_{k=0}^{\infty}\langle f\rangle_{Q_{k+1}(x)}-\langle f\rangle_{Q_{k}(x)}\Big{|}_{Y}\leq\Big{(}\sum_{k=0}^{M-K}+\sum_{k=M-K+1}^{\infty}\Big{)}|\langle f\rangle_{Q_{k+1}(x)}-\langle f\rangle_{Q_{k}(x)}|_{Y}$
$\lesssim_{n}\varepsilon\sum_{k=0}^{M-K}\omega(\ell(Q_{k}(x)))+\sup_{\ell(Q)\leq 2^{K}}\mathcal{O}^{\omega}(f;Q)\sum_{k=M-K+1}^{\infty}\omega(\ell(Q_{k}(x)))$
$\lesssim_{C_{\operatorname{db}}}\varepsilon\int_{2^{K}}^{2^{M+1}}\frac{\omega(t)}{t}\,\mathrm{d}t+\sup_{\ell(Q)\leq 2^{K}}\mathcal{O}^{\omega}(f;Q)\int_{0}^{2^{K}}\frac{\omega(t)}{t}\,\mathrm{d}t$
$\lesssim_{C_{\operatorname{db}}}\Big{[}\varepsilon\Big{(}\frac{1}{\omega(2^{M+1})}\int_{2^{K}}^{2^{M+1}}\frac{\omega(t)}{t}\,\mathrm{d}t\Big{)}$
$\qquad\qquad\qquad+\frac{\omega(2^{K})}{\omega(2^{M+1})}\sup_{\ell(Q)\leq 2^{K}}\mathcal{O}^{\omega}(f;Q)\Big{(}\frac{1}{\omega(2^{K})}\int_{0}^{2^{K}}\frac{\omega(t)}{t}\,\mathrm{d}t\Big{)}\Big{]}\omega(\ell(Q_{0}))$
$\lesssim\Big{[}\varepsilon[\omega]_{*}+\frac{\omega(2^{K})}{\omega(2^{N+1})}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}[\omega]_{*}\Big{]}\omega(|x-y|).$
By the condition $\omega(\infty)=\infty$, we choose $N$ large enough and bound the bracketed term from above by $\leq 2\varepsilon[\omega]_{*}.$ Repeating the same proof with $y$ in place of $x$, we control the right sum on the line (29) and thus obtain $|f(x)-f(y)|\leq 4\varepsilon[\omega]_{*}\omega(|x-y|),$ that is, $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$
Then, we show that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$ Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y)$, $\varepsilon>0,$ and $R>0$ be such that $\operatorname{osc}^{\omega}_{\delta}(f)\leq\varepsilon,$ if $\delta\geq R.$ Let $M\gg R$ be a large constant, which we will specify later, and $Q\in\mathcal{Q}$ a cube such that $\ell(Q)\geq M.$ For every $y\in Q,$ denote by $Q_{R}(y)$ the cube centered at $y$ and of $\operatorname{diam}(Q_{R}(y))=R.$ Then, we have
$\fint_{Q}\int_{Q}|f(x)-f(y)|_{Y}\,\mathrm{d}x\,\mathrm{d}y=\fint_{Q}\Big{(}\int_{Q\setminus Q_{R}(y)}+\int_{Q\cap Q_{R}(y)}\Big{)}|f(x)-f(y)|_{Y}\,\mathrm{d}x\,\mathrm{d}y$
$\leq\varepsilon\fint_{Q}\int_{Q\setminus Q_{R}(y)}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y+\sup_{\delta\leq R}\operatorname{osc}^{\omega}_{\delta}(f)\fint_{Q}\int_{Q\cap Q_{R}(y)}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y$
$\leq\Big{[}\varepsilon\fint_{Q}\fint_{Q}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y+\sup_{\delta\leq R}\operatorname{osc}^{\omega}_{\delta}(f)\fint_{Q}\frac{1}{|Q|}\int_{Q_{R}(y)}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y\Big{]}|Q|.$
By repeating the bound (28), the left term in the square bracket is bounded from above by $\lesssim_{n,C_{\operatorname{db}}}\varepsilon\omega(\ell(Q)),$ which is a bound of the correct form. For the right term we bound
$\lesssim_{n}\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\left(\frac{R}{\ell(Q)}\right)^{n}\omega(R)\leq\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\omega(R)\leq\varepsilon\omega(\ell(Q));$
as $\omega(\infty)=\infty,$ and $M$ is chosen sufficiently large. ∎
3. Approximation in infinite dimensional spaces
In this section we study smooth approximations in Banach spaces with respect to the $\dot{C}^{0,\omega}$ seminorm. In Subsections 3.1 and 3.2 we establish results for real-valued functions $f:X\to\mathbb{R}$, where the approximating functions are not only $C^{k}$ or $C^{1,\alpha}$ smooth but also Lipschitz and with bounded support. In the last Subsection 3.3 we obtain $C^{\infty}$ smooth approximations of Banach-valued mappings $f:c_{0}(\mathcal{A})\to Y,$ that are not necessarily Lipschitz, but have bounded support and belong to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ As a corollary, we deduce $C^{\infty}$ and $\dot{\operatorname{VC}}^{0,\omega}$ and approximations with bounded support for all $\dot{\operatorname{VC}}^{0,\omega}$ functions defined on $\mathbb{R}^{n}$ and with values in any Banach space.
As we mentioned in the introduction, for general background on smooth analysis on Banach spaces, including smooth renomings and approximations, we refer the reader to the monographs [BenyaminiLindenstrauss] by Benyamini and Lindenstrauss; [DevilleGodefroyZizler] by Deville, Godefroy, and Zizler; and [HajJoh10] by Hájek and Johanis.
Assumptions on the modulus
Throughout this whole Section 3 we assume that the modulus satisfies all the assumptions of Theorem 1.2, i.e. $\omega$ is non-decreasing and satisfies (2), (3) and (4).
Notation and basic definitions
By $\operatorname{Lip}(X)$ we denote the class of real-valued Lipschitz functions (not necessarily bounded) on $X.$ We say that $g:X\to\mathbb{R}$ is $L$-Lipschitz provided that $|g(x)-g(y)|\leq L\|x-y\|$ for every $x,y\in X,$ and we denote the minimal Lipschitz constant of $g$ by $\operatorname{Lip}(g).$
We will denote by $X^{*}$ the (continuous) dual of $X,$ and the dual norm in $X^{*}$ by $\|\cdot\|_{*},$ that is,
$\|\Lambda\|_{*}:=\sup\{|\Lambda(v)|\,:\,\|v\|\leq 1\},$
for every $\Lambda\in X^{*}.$
When speaking of differentiability of functions or mappings $f:X\to Y$, we always mean differentiability in the Fréchet sense. We say, of course, that $f$ is of class $C^{k}(X,Y),$ when $f$ has Fréchet derivatives up to order $k,$ and those derivatives are continuous on $X.$
[…]
For any class of functions $\mathcal{C}(X,Y),$ the subscripted set $\mathcal{C}_{\operatorname{bs}}(X,Y)$ consists of functions $h:X\to Y$ of class $\mathcal{C}$ that have bounded support, meaning that there exists $R>0$ so that $h(x)=0$ for every $x\in X\setminus B(0,R).$
Finally, a bump function $h:X\to Y$ is a non-zero function with bounded support.
3.1. Lipschitz and smooth approximations
In a real normed space $X$, we consider real-valued functions ``$f"$ of the class $\dot{\operatorname{VC}}^{0,\omega}(X,\mathbb{R})$, the last abbreviated by $\dot{\operatorname{VC}}^{0,\omega}(X).$ By Theorem 1.2 these functions ``$f"$ can be approximated by functions, say ``$g",$ of class $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ with bounded support, in the $\dot{C}^{0,\omega}(X)$ seminorm. Our goal is to further approximate these functions ``$g"$ by Lipschitz functions (whose regularity will vary depending on the smoothness properties of $X$) with bounded support. In approximating ``$g"$ a crucial step is to reduce the problem to the problem of approximating Lipschitz functions by smooth Lipschitz functions with good control on the Lipschitz constants. The main results of this subsection are Theorem 3.3, valid for arbitrary normed spaces, and Theorem 3.5, valid for separable normed spaces $X$ with a fixed degree of smoothness. The proofs of these two Theorems 3.3 and 3.5 employ several technical lemma, which will be reused in the later Subsections 3.2 and 3.3.
We begin by proving these technical lemmas and for this purpose, it will be useful to consider the minimal modulus of continuity of a function $f:X\to\mathbb{R},$ i.e.,
$\omega_{f}(t):=\sup\{|f(x)-f(y)|\,:\,\|x-y\|\leq t\},\quad t>0.$
Note that $\omega_{f}:[0,\infty]\to[0,\infty]$ is non-decreasing, and if $f$ is uniformly continuous, then $\lim_{t\to 0}\omega_{f}(t)=0.$ Moreover, if $f\in\dot{C}^{0,\omega}(X),$ then $\omega_{f}\leq\|f\|_{\dot{C}^{0,\omega}(X)}\omega,$ and especially $\omega_{f}(t)$ is finite at every $t,$ but, in general, $\omega$ and $\omega_{f}$ are incomparable. Furthermore, since $X$ is a normed space, it is easy to see that $\omega_{f}$ is sub-additive, i.e. $\omega_{f}(s+t)\leq\omega_{f}(t)+\omega_{f}(s)$ for all $t,s>0.$ This implies
$\omega_{f}(t)\leq 2t\,\omega_{f}(1),\quad\text{for all}\quad t\geq 1.$ (30)
Lemma 3.1.
Let $X$ be a normed space and $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X).$ Then, there exists a sequence $(f_{n})_{n}\subset\operatorname{Lip}(X)$ of functions converging to $f$ in the $\dot{C}^{0,\omega}(X)$-seminorm. Moreover, if $f$ has bounded support, the sequence can be taken $(f_{n})_{n}\subset\operatorname{Lip}_{\operatorname{bs}}(X)$ to have bounded supports.
Proof.
As $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$ by the comments preceding the present lemma, the minimal modulus of continuity $\omega_{f}$ is sub-additive and satisfies $\omega_{f}(t)\to 0$, as $t\to 0;$ then, it is a known result that the sequence
$f_{n}(x)=\inf_{y\in X}\{f(y)+n\|x-y\|\},\quad x\in X,$
converges uniformly to $f$ in $X$ and each $f_{n}$ is $n$-Lipschitz; for a proof see e.g. [HajJoh14, p. 408].
Now, let us see that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\to 0,$ as $n\to\infty.$ Indeed, given $\varepsilon>0$ there exists $\delta>0$ so that $\|x-z\|\leq\delta$ implies $|f(x)-f(z)|\leq\varepsilon\omega(\|x-z\|).$ Let us take $N\in\mathbb{N}$ large enough so that $\sup_{X}|f-f_{n}|\leq\varepsilon\omega(\delta)$ for each $n\geq N.$ Using that $\omega$ is non-decreasing, this gives the estimate
$|(f-f_{n})(x)-(f-f_{n})(z)|\leq 2\|f-f_{n}\|_{\infty}\leq\varepsilon\omega(\delta)$ $\leq\varepsilon\omega(\|x-z\|),$
$\text{whenever}\quad\|x-z\|\geq\delta,\,n\geq N.$
Now, for fixed $x,z\in X$ such that $\|x-z\|\leq\delta,$ we proceed as follows. Given $\eta>0,$ let $y=y(x,n,f,\eta)$ be so that
$f_{n}(x)\geq f(y)+n\|x-y\|-\eta.$
This choice of $y$ and the definition of $f_{n}$ yield
$(f-f_{n})(x)$ $-(f-f_{n})(z)\leq f(x)-f(y)-n\|x-y\|+\eta-f(z)+f_{n}(z)$
$\leq f(x)-f(y)-n\|x-y\|+\eta-f(z)+f(y+z-x)+n\|z-(y+z-x)\|$
$\leq f(x)-f(z)-(f(y)-f(y+z-x))+\eta\leq\varepsilon\omega(\|x-z\|)+\varepsilon\omega(\|x-z\|)+\eta.$
Letting $\eta\to 0,$ we get
$(f-f_{n})(x)-(f-f_{n})(z)\leq 2\varepsilon\omega(\|x-z\|).$
The same argument swapping $x$ and $z,$gives $|(f-f_{n})(x)-(f-f_{n})(z)|\leq 2\varepsilon\omega(\|x-z\|),$ for every $n\in\mathbb{N}.$ We conclude that $\|f-f_{n}\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon,$ for $n\geq N.$
For the assertion concerning the boundedness of the supports, we may assume that $f(y)=0,$ whenever $\|y\|\geq R,$ for some $R>0.$ We will localize the infimum defining $f_{n}.$ For each $y\in X,$ it follows from the definition of $f_{n}$, and the fact that $f\in\dot{C}^{0,\omega_{f}}(X),$ that
$f(y)+n\|x-y\|$ $\geq f(x)-\|f\|_{\dot{C}^{0,\omega_{f}}(X)}\omega_{f}(\|x-y\|)+n\|x-y\|$
$\qquad\geq f_{n}(x)-\|f\|_{\dot{C}^{0,\omega_{f}}(X)}\omega_{f}(\|x-y\|)+n\|x-y\|.$ (31)
[…]
$f_{n}(x)=\inf_{y\in B(x,1)}\{f(y)+n\|x-y\|\},\quad x\in X.$
Now, if $\|x\|\geq R+1,$ then $\|y\|\geq R$, and so $f(y)=0,$ whenever $y\in B(x,1).$ Thus, by the previous formula, we deduce $f_{n}(x)=\inf_{y\in B(x,1)}\{n\|x-y\|\}=0,$ which shows that $f_{n}$ has bounded support, for $n$ large enough.
In the following lemma we show that a uniform approximation by Lipschitz functions with controlled Lipschitz constants yields a $\dot{C}^{0,\omega}(X)$ approximation.
Lemma 3.2.
Let $f:X\to\mathbb{R}$ and $(f_{n})_{n}\subset\operatorname{Lip}(X)$ be a sequence such that $\limsup_{n}\mathrm{Lip}(f_{n})<\infty$ and $f_{n}$ converges to $f$ uniformly on $X.$ Then, $\limsup_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}(X)}=0.$
Proof.
By passing to a subsequence, we may assume that $L:=\sup_{n}\mathrm{Lip}(f_{n})<\infty$. Then, $f$ is Lipschitz on $X,$ with $\operatorname{Lip}(f)\leq L.$ Given $\varepsilon>0,$ by $\lim_{t\to 0}t/\omega(t)=0,$ we find $\delta>0$ so that
$\frac{t}{\omega(t)}\leq\frac{\varepsilon}{1+L},\quad\text{whenever}\quad t\leq\delta.$
Let $N\in\mathbb{N}$ be such that $\sup_{X}|f_{n}-f|\leq\varepsilon\omega(\delta)$ for all $n\geq N.$ Because $f$ and each $f_{n}$ is $L$-Lipschitz, for any two distinct points $x,z\in X$ with $\|x-z\|\leq\delta,$ we have
$\max\left\{\frac{|f(x)-f(z)|}{\omega(\|x-z\|)},\frac{|f_{n}(x)-f_{n}(z)|}{\omega(\|x-z\|)}\right\}\leq L\frac{\|x-z\|}{\omega(\|x-z\|)}\leq\varepsilon;$
then, triangle inequality gives
$\frac{|(f-f_{n})(x)-(f-f_{n})(z)|}{\omega(\|x-z\|)}\leq\frac{|f(x)-f(z)|}{\omega(\|x-z\|)}+\frac{|f_{n}(x)-f_{n}(z)|}{\omega(\|x-z\|)}\leq 2\varepsilon.$
And for scales $\|x-z\|\geq\delta,$ observe that for $n\geq N,$ using that $\omega$ is non-decreasing, we have
$|(f-f_{n})(x)-(f-f_{n})(z)|\leq 2\sup_{X}|f-f_{n}|\leq 2\varepsilon\omega(\delta)\leq 2\varepsilon\omega(\|x-z\|).$
We are now ready to prove our first main Theorem 3.3 on Lipschitz approximations. It does not involve smoothness, but holds in any normed space and will be a key ingredient in the proofs of Theorems 3.10 and 3.12 of later sections.
Theorem 3.3 (Lipschitz approximation of $\dot{\operatorname{VC}}^{0,\omega}$).
Let $X$ be a normed space. Then, the following hold:
  • (i)
    $\overline{\mathrm{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$
  • (ii)
    $\overline{\mathrm{Lip}_{\operatorname{bs}}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}(X).$
Proof.
The inclusion ``"``"`$\subset$"` ` ⊂ " follows immediately from Lemma 3.1. For the converse inclusion ``","```$\supset$",` ` ⊃ " , we observe that condition (3) gives $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$, and then recall that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ is closed.
By Theorem 1.2 we know that if $f\in\dot{\operatorname{VC}}^{0,\omega}(X),$ and $\varepsilon>0,$ we can find $g\in\dot{\operatorname{VC}}^{0,\omega}(X)$ with bounded support and $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Then, Lemma 3.1 provides us with $h\in\operatorname{Lip}_{\operatorname{bs}}(X)$ so that $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and therefore $\|f-h\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon;$ this shows the inclusion ``"."```$\supset$".` ` ⊃ " .
For the converse implication ``"``"`$\subset$"` ` ⊂ ", it is enough to observe that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X);$ see Remark 2.1. ∎
Performing smooth approximations in the appropriate Banach spaces will require more work. We first show that if a normed space $X$ has a $C^{k}$ and Lipschitz bump function, from now on a $C^{k}\cap\operatorname{Lip}$ bump, then a $C^{k}\cap\operatorname{Lip}$ approximation with uniformly bounded Lipschitz constants, of a Lipschitz function $g$ with bounded support, can be upgraded to a $C^{k}\cap\operatorname{Lip}$ approximation with uniformly bounded Lipschitz constants and bounded supports.
Lemma 3.4.
Let $k\in\mathbb{N}\cup\{\infty\}$ be fixed and let $X$ be a normed space that admits a $C^{k}\cap\operatorname{Lip}$ bump. Let $g\in\operatorname{Lip}(X)$ and $(g_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ be a sequence of functions converging uniformly to $g$ on $X,$ and with $\limsup_{n}\operatorname{Lip}(g_{n})<\infty.$
Then, if $g$ has bounded support, there exists a sequence $(h_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ with bounded supports that converge to $g$ uniformly (on $X$) and moreover $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$
Proof.
Let $g$ be as in the assumption, and let $R>0$ be so that $g=0$ outside the ball $B(0,R).$ We first construct a suitable bump that decays from one to zero over the annular region $B(0,\lambda R)\setminus B(0,2R),$ for a certain $\lambda>2$ to be fixed later. Since $X$ admits a $C^{k}\cap\operatorname{Lip}$ bump, by [DevilleGodefroyZizler, Proposition II, 5.1] there exist a constant $10<a<$ 1 and a function $\psi:X\to[0,\infty)$ so that $\psi\in\operatorname{Lip}(X)\cap C^{k}(X\setminus\{0\})$ and
$a\|x\|\leq\psi(x)\leq\|x\|,\quad x\in X.$ (32)
The statement and proof of [DevilleGodefroyZizler, Proposition II, 5.1] are written for $k=1,$ but they clearly hold true for $k\in\mathbb{N}\cup\{\infty\}$ as well. Now, pick a function $\theta:\mathbb{R}\to[0,1]$ so that $\theta\in C^{\infty}(\mathbb{R})\cap\operatorname{Lip}(\mathbb{R}),$ and $\theta(t)=1,$ whenever $t\leq 2R$, and $\theta(t)=0,$ whenever $t\geq 4R.$ Then, define a bump function by
$\varphi:X\to\mathbb{R},\qquad\varphi(x)=\theta(\psi(x)),\quad x\in X.$ (33)
By the properties of $\theta$ and $\psi,$ it is immediate that $\varphi\in C^{k}(X)\cap\operatorname{Lip}(X)$, and that $\varphi$ takes values in $[0,1].$ Also, if $\|x\|\leq 2R,$ one has $\psi(x)\leq 2R$ by (32), and thus $\varphi(x)=\theta(\psi(x))=1.$ Similary, we deduce that $\varphi(x)=\theta(\psi(x))=0$ for those $x\in X$ such that $\|x\|\geq 4R/a=:\lambda.$
Now, by assumption we can find a sequence $(g_{n})_{n}\subset C^{k}(X)\cap\operatorname{Lip}(X)$ converging uniformly to $g$ on $X,$ and, after passing to a subsequence, with the property $\sup_{n}\operatorname{Lip}(g_{n})<\infty.$ Define $h_{n}:=\varphi g_{n}$ for each $n\in\mathbb{N},$ where $\varphi$ is as in (33). Since $\varphi$ vanishes outside the ball $B(0,4R/a),$ the function $g_{n}$ has bounded support. Also,
$|h_{n}(x)-g(x)|=|\varphi(x)g_{n}(x)-g(x)|\leq|\varphi(x)-1||g_{n}(x)|+|g_{n}(x)-g(x)|.$ (34)
The second term in the sum converges to 00 uniformly on $x.$ For the first term, note that if $\|x\|\leq 2R,$ then $\varphi(x)=1,$ and so $|\varphi(x)-1||g_{n}(x)|=0.$ When $\|x\|\geq 2R,$ we have $\lim_{n}|g_{n}(x)|=|g(x)|=0,$ uniformly on those $x,$ by the bounded support of $g.$ Since also $\varphi$ takes values in $[0,1],$ we conclude that the first term tends uniformly to 00 as $n\to\infty.$ Now we have shown that
$\lim_{n\to\infty}\sup_{x\in X}|h_{n}(x)-g(x)|\to 0.$
Concerning the regularity of $h_{n},$ note that obviously $h_{n}\in C^{k}(X)$ because both $\varphi$ and $g_{n}$ are of class $C^{k}.$ To verify that $\limsup_{n}\operatorname{Lip}(h_{n})<\infty,$ we estimate the derivative of $h_{n}$ by
$\|Dh_{n}(x)\|_{*}\leq\varphi(x)\|Dg_{n}(x)\|_{*}+|g_{n}(x)|\|D\varphi(x)\|_{*}\leq\operatorname{Lip}(g_{n})+\operatorname{Lip}(\varphi)(\sup_{X}|g_{n}-g|+\sup_{X}|g|).$
Then $\operatorname{Lip}(g_{n})\leq\sup_{m}\operatorname{Lip}(g_{m}),$ $\sup_{X}|g_{n}-g|\to 0$, and $\sup_{X}|g|<\infty$ because $g$ is Lipschitz with bounded support. This shows $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$
Now, we combine Lemmas 3.1, 3.2, Theorem 3.3, and Lemma 3.4, with an approximation result for separable spaces that admit a $C^{k}\cap\operatorname{Lip}$ bump, and obtain our most general theorem concerning approximation of $\dot{\operatorname{VC}}^{0,\omega}$ functions by $C^{k}\cap\operatorname{Lip}$ functions in separable spaces.
Theorem 3.5 (Smooth approximation in separable spaces).
Let $k\in\mathbb{N}\cup\{\infty\}$ and let $X$ be a separable normed space admitting a $C^{k}\cap\operatorname{Lip}$ bump. Then, the following hold:
  • (i)
    $\overline{C^{k}(X)\cap\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X),$
  • (ii)
    $\overline{C^{k}_{\operatorname{bs}}(X)\cap\mathrm{Lip}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}(X).$
Proof.
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\varepsilon>0.$ By Lemma 3.1 we find an approximation $g\in\operatorname{Lip}(X)$ so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Now, since $g:X\to\mathbb{R}$ is a Lipschitz function, and $X$ has a $C^{k}\cap\operatorname{Lip}$ bump, as a consequence of a result of Hájek and Johanis [HajJoh10, Corollary 15] (here we use that $X$ is separable), we find a sequence $(g_{n})_{n}$ of $C^{k}\cap\operatorname{Lip}$ functions converging uniformly to $g$ such that $\operatorname{Lip}(g_{n})\leq\lambda\operatorname{Lip}(g),$ for all $n,$ and for some absolute constant $\lambda>0$ that may depend on the space $X.$ Since obviously the function $g$ and the sequence $(g_{n})_{n}$ satisfy all the assumptions of Lemma 3.2, and hence $\lim_{n}\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}=0,$ completing the proof of the inclusion ``"."```$\supset$".` ` ⊃ " . For the reverse inclusion recall that the assumption $t/\omega(t)\to 0,$ as $t\to 0,$ guarantees that $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X);$ and $\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X)$ is a closed subspace of ${\dot{C}^{0,\omega}(X)}.$
Assume that $f\in\dot{\operatorname{VC}}^{0,\omega}(X)$ and let $\varepsilon>0.$ According to Theorem 1.2 we find $g\in\operatorname{Lip}(X)$ with bounded support so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Now, let $(g_{n})_{n}$ be the sequence we used in part (i) (but associated with this new $g$), and then Lemma 3.4 guarantees a $C^{k}\cap\operatorname{Lip}$ approximation with bounded support $(h_{n})_{n}$ so that $\lim_{n}\sup_{X}|h_{n}-g|=0$ and $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$ Now, we apply Lemma 3.2 to deduce that $\lim_{n}\|h_{n}-g\|_{\dot{C}^{0,\omega}(X)}=0$ as well. Thus we pick some $h\in(h_{n})_{n}\subset C^{k}(X)\cap\mathrm{Lip}_{\operatorname{bs}}(X)$ such that $\|g-h\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ which in turn implies
$\|f-h\|_{\dot{C}^{0,\omega}(X)}\leq\|f-g\|_{\dot{C}^{0,\omega}(X)}+\|g-h\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon.$
Thus we have shown the inclusion ``"."```$\supset$".` ` ⊃ " . For the converse inclusion ``"``"`$\subset$"` ` ⊂ ", it is enough to observe again (as we did at the end of the proof of Theorem 3.3) that $\mathrm{Lip}_{\operatorname{bs}}(X)\subset\dot{\operatorname{VC}}^{0,\omega}(X)$ and that $\dot{\operatorname{VC}}^{0,\omega}(X)$ is a closed subspace of ${\dot{C}^{0,\omega}(X)},$ again, by Remark 2.1.
Remark 3.6.
Clarifications and remarks concerning Theorem 3.5 and its proof.
  1. (1)
[…]
  2. (2)
    Let us now recall the smoothness of the canonical, or equivalent, norms of some classical separable spaces. The canonical norms of the spaces $X=\ell_{p}$ or $L^{p},$ for $1<p<\infty$, have the following order of smoothness $k=k(p):$
    $k(p)=\begin{cases}\infty\qquad&\text{ if }p\text{ is even},\\ p-1,\qquad&\text{ if }p\text{ is odd},\\ \left\lfloor p\right\rfloor,\qquad&\text{ if }p\text{ is not an integer}.\end{cases}$
    Also, for a compact metric space $K$ which is scattered, let $C(K)$ be the Banach space consisting of real-valued continuous functions on $K$, equipped with the supremum norm. Then $C(K)$ admits an equivalent $C^{\infty}$-norm. Recall that a set $S$ is scattered if every nonempty subset of $S$ contains a (relatively) isolated point. We refer the reader to [DevilleGodefroyZizler, Theorems V.1.1 and V.1.8] for detailed proofs and statements of these theorems. Then, according to point (1) above, Theorem 3.5 applies for $X=L^{p},\,\ell_{p}$ with $k_{p}$ as above, and for $X=C(K)$ with $K$ countable and $k=\infty.$
  3. (3)
    For $X=c_{0},$ see Section 3.3 below, Theorem 3.5 also holds for $k=\infty,$ but for this particular $X,$ we will obtain much more in Theorem 3.12 below.
  4. (4)
    In the proof of Theorem 3.5 we employed [HajJoh10, Corollary 15] for our real-valued functions, but this result holds for target spaces $Y=B_{0}(V)$ (real-valued functions $f$ in a topological space $V$ with $f(v)\to 0$ as $v\to v_{0}$, for some fixed $v_{0}\in V$); or $Y=C_{u}(P)$ (uniformly continuous bounded functions on a metric space $P$). Naturally, the particular case $Y=\mathbb{R}$ is covered by these spaces, e.g., when $Y=C_{u}(P)$ for $P=\{0\}\subset\mathbb{R}$.
Now, as a corollary of Theorem 3.5, we provide $C^{\infty}\cap\operatorname{Lip}$ approximation of $\dot{\operatorname{VC}}^{0,\omega}(X)$ functions for separable Hilbert spaces $X$. In particular, notice that this extends Theorem 1.6 to infinite dimensions.
Corollary 3.7.
Let $X$ be a separable Hilbert space. Then, the following hold:
  • (i)
    $\overline{C^{\infty}(X)\cap\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X),$
  • (ii)
    $\overline{C^{\infty}_{\operatorname{bs}}(X)\cap\mathrm{Lip}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}(X).$
Proof.
Both statements (i) and (ii) follow immediately from Theorem 3.5, since a Hilbert space always admits a $C^{\infty}\cap\operatorname{Lip}$ bump. For the sake of completeness, we exhibit an elementary construction of such a bump.
The function $\psi:X\to\mathbb{R}$ given by $x\mapsto\psi(x)=\|x\|^{2}$ is of class $C^{\infty}(X)$ with the Fréchet derivative $D\psi(x):X\to\mathbb{R},$ at every $x\in X$, given by $D\psi(x)(v)=\langle 2x,v\rangle$ for all $v\in X.$ Take a function $\theta:\mathbb{R}\to[0,1]$ of class $C^{\infty}(\mathbb{R})\cap\operatorname{Lip}(\mathbb{R})$ with $\theta(t)=1$ for all $t\leq 1/2$ and $\theta(t)=0$ for all $t\geq 1.$ Then, the map $\varphi=\theta\circ\psi:X\to[0,1]$ is of class $C^{\infty}(X)$ with $\varphi(x)=1$ for all $\|x\|\leq 1/2$ and $\varphi(x)=0$ for all $\|x\|\geq 1.$ Also, the Fréchet derivative of $\varphi$ satisfies
$\|D\varphi(x)\|_{*}=\|\theta^{\prime}(\psi(x))\mathcal{X}_{B(0,1)}(x)2x\|_{*}\leq 2\operatorname{Lip}(\theta),$
showing that $\varphi$ is Lipschitz on $X.$ Therefore, $\varphi$ is a $C^{\infty}\cap\operatorname{Lip}$ bump on $X.$
3.2. Approximation in super-reflexive spaces
[…]
$\|x+h\|^{1+\alpha}+\|x-h\|^{1+\alpha}-2\|x\|^{1+\alpha}\leq C\|h\|^{1+\alpha},\quad\text{for all}\quad x,h\in X.$ (35)
See also [BenyaminiLindenstrauss, pp. 412–413] for a proof of this equivalence. This property is often rephrased by saying that $X$ admits a renorming with modulus of smoothness of power type $1+\alpha.$ Since the function $x\mapsto\psi(x)=\|x\|^{1+\alpha}$ is convex and continuous on $X,$ then (35) implies that $\psi$ is of class $C^{1,\alpha}(X);$ and with $\|D\psi\|_{\dot{C}^{0,\alpha}(X,X^{*})}\lesssim_{\alpha,C}1$; see [DevilleGodefroyZizler, Lemma V.3.5].
We remind that a function $f:X\to\mathbb{R}$ belongs to the class $C^{1,\alpha}(X)$ if $f$ is Fréchet differentiable at every point $x\in X$ and the Fréchet derivative $Df:X\to X^{*}$ is $\alpha$-Hölder continuous on $X,$ namely, that
$\|Df\|_{\dot{C}^{0,\alpha}(X,X^{*})}:=\sup\left\{\frac{\|Df(x)-Df(y)\|_{*}}{\|x-y\|^{\alpha}}\,:\,x,y\in X,\,x\neq y\right\}<\infty.$
In Hilbert spaces the Lasry–Lions regularization theorem (see [LasryLions] or [HajJoh10, p. 408]) provides uniform approximation of Lipschitz functions by $C^{1,1}\cap\operatorname{Lip}$ functions. Cepedello-Boiso [Cepedello, Theorem 1] used an ingenious variant of the Lasry-Lions technique to obtain uniform approximation of Lipschitz functions in super-reflexive spaces by $C^{1}$ functions whose derivatives are $\alpha$-Hölder on bounded subsets, but not globally on $X.$ These approximations are, of course, Lipschitz on bounded sets, but not globally Lipschitz. However, for our purposes, we need to approximate Lipschitz functions uniformly by globally $C^{1,\alpha}\cap\operatorname{Lip}$ functions, and we need a good control on the Lipschitz constants of the approximations. In Theorem 3.8(2) below, we have obtained uniform approximations that are globally $C^{1,\alpha}$ and globally Lipschitz on $X$, preserving the Lipschitz constants up to a multiplicative factor depending only on $\alpha$ and $X.$ In the proof, we use a recent result on $C^{1,\omega}$ extensions for jets [AzaMud21] to construct $C^{1,\alpha}$ bump functions with all the properties stated in part (1) of Theorem 3.8. This is combined with a method of Hájek and Johanis [HajJoh10, Proposition 1] to glue up a suitable sequence of these bumps. Both parts (1) and (2) in Theorem 3.8 are interesting to the theory of smooth approximation, and also will be essential for the approximation of $\dot{\operatorname{VC}}^{0,\omega}$ functions.
Let us mention that, after making the first version of this article public, it has come to our attention that part (2) of Theorem 3.8 was very recently discovered by Johanis [Jo24] too, by means of a different proof. A benefit of Johanis’ proof is that it provides the sharp constant $\kappa=1.$ Since we only need an absolute control on $\kappa$ for our purposes, and since we consider part (1) of independent interest as well, we have chosen to include our original proof of Theorem 3.8(2).
Theorem 3.8.
[…]
  1. (1)
    For every set $S\subset X$ there exists $h_{S}:X\to[0,1]$ of class $C^{1,\alpha}(X)\cap\operatorname{Lip}(X)$ so that:
    $h_{S}(x)=0,$ for every $x\in S$;
    $h_{S}(x)=1,$ whenever $d(x,S)\geq 1;$
    $Dh_{S}(x)=0,$ for every $x\in S\cup\{y\in X\,:\,d(y,S)\geq 1\};$ and
    $\operatorname{Lip}(h_{S})+\|Dh_{S}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa.$
  2. (2)
    Given an $L$-Lipschitz function $g:X\to\mathbb{R}$ and $\varepsilon>0,$ there exists $h\in C^{1,\alpha}(X)$ so that $h$ is $\kappa L$-Lipschitz, and $\sup_{X}|g-h|\leq\varepsilon.$
Proof of Theorem 3.8.
In the cases $S=\varnothing$ or $\{\operatorname{\mathit{d}}(\cdot,S)\geq 1\}=\varnothing,$ we simply take $h_{S}\equiv 1$ in the first case, and $h_{S}\equiv 0$ in the latter. Assume from now on that both sets are nonempty. On the set
$E:=S\cup\{x\in X\,:\,d(x,S)\geq 1\},$
we define a $1$-jet $(f,G):E\to\mathbb{R}\times X^{*}$ by setting
$f(x)=0,\quad\text{if}\quad x\in S;\qquad f(x)=1,\quad\text{if}\quad d(x,S)\geq 1;\qquad G(x)=0,\quad\text{for all}\quad x\in E.$
By separating into cases, it is immediate to verify the existence of $M=M(\alpha)>0$ for which
$f(y)+G(y)(x-y)-f(z)-G(z)(x-z)\leq\frac{M}{1+\alpha}\left(\|x-y\|^{1+\alpha}+\|x-z\|^{1+\alpha}\right),$
for all $y,z\in E,$ $x\in X;$ indeed, $G(y)=G(z)=0$ is the zero functional, and for $f(y)-f(z),$ by symmetry, it is enough to verify the case $y\in S$ and $\operatorname{\mathit{d}}(z,S)\geq 1$ and $x\in X$ arbitrary, but this case is clear by inspection, since
$f(y)-f(z)=1\leq\|y-z\|\leq\|y-z\|^{1+\alpha}\leq\frac{M}{1+\alpha}\Big{(}\|x-y\|^{1+\alpha}+\|x-z\|^{1+\alpha}\Big{)}.$
Thus, applying [AzaMud21, Theorem 4.1], there exists a constant $\kappa_{0}=\kappa_{0}(\alpha,C)$, and a function $F:X\to\mathbb{R}$ of class $C^{1,\alpha},$ bounded and Lipschitz, with
$\|F\|_{\infty}+\operatorname{Lip}(F)+\|DF\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa_{0}M=:\kappa_{1},$ (36)
and so that $F=f$ and $DF=G=0$ on $E.$ Although [AzaMud21, Theorem 4.1] is formulated for Hilbert spaces, it generalizes immediately to super-reflexive Banach spaces satisfying property (35), as pointed out in [AzaMud21, Remark 4.7]; see also [AzaMud21, Theorem 1.9]. Let us examine the extension formula $F$, since we need some of its properties. At the same time, for the reader’s convenience, we sketch some of the key steps in the proof of [AzaMud21, Theorem 4.1]. With $M$ as above,
$M^{*}:=\max\Big{\{}3\big{(}\|f\|_{\infty}+\|G\|_{\infty}\big{)},\,\frac{M}{1+\alpha}\Big{\}},$
and the functions $m,g:X\to\mathbb{R}$ by
$m(x)$ $:=\max\Big{\{}-2\Big{(}\|f\|_{\infty}+\|G\|_{\infty}\Big{)},\,\sup_{z\in E}\Big{\{}f(z)+G(z)(x-z)-M^{*}\|x-z\|^{1+\alpha}\Big{\}}\Big{\}},$
$g(x)$ $:=\min\Big{\{}2\Big{(}\|f\|_{\infty}+\|G\|_{\infty}\Big{)},\,\inf_{y\in E}\Big{\{}f(y)+G(y)(x-y)+M^{*}\|x-y\|^{1+\alpha}\Big{\}}\Big{\}}.$
For a suitable number $A$ (depending only on $\alpha$ and $C$), $F$ is defined to be the $AM^{*}t^{1+\alpha}$-strongly paraconvex envelope of $g$, that is,
$F(x):=\sup\{h(x)\,:\,h\text{ is }AM^{*}t^{1+\alpha}\text{-strongly paraconvex,and }h\leq g\text{ on }X\},\quad x\in X.$
As defined in [AzaMud21], a function $h:\psi\to\mathbb{R}$ is $Lt^{1+\alpha}$-strongly paraconvex, for $L>0,$ if
$\psi(\lambda u+(1-\lambda)v)-\lambda\psi(u)-(1-\lambda)\psi(v)\leq\lambda(1-\lambda)L\|u-v\|^{1+\alpha},\quad u,v\in X,\,\lambda\in[0,1].$
By the comment just after formula (35), the function $u\mapsto\psi(u):=-\|u\|^{1+\alpha}$ is $C^{1,\alpha}$, and their Fréchet derivative $D\psi$ satisfies $\|D\psi\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq B(\alpha,C).$ Thus, using the Fundamental Theorem of Calculus and some elementary computations, $\psi$ is $At^{1+\alpha}$-strongly paraconvex for some $A=A(\alpha,C);$ see the argument near the end of [AzaMud21, the proof of Lemma 3.6]. It turns out that then both $m$ and $(-g)$ are $AM^{*}t^{\alpha}$-strongly paraconvex. Also, $m\leq g$ on $X,$ and these properties permit to prove the regularity $F\in C^{1,\alpha}(X)$, along with (36) and that $(F,DF)=(f,G)$ on $E.$ See [AzaMud21, Theorems 1.9 and 4.1] for further explanations and details of these concepts and their proofs. Now, by the definition of $F,$ the pointwise relation
$m(x)\leq F(x)\leq g(x),\quad x\in X,$
holds true. Since $G\equiv 0$ and $\|f\|_{\infty}=1,$ the definition of $m$ and $g$ and this estimate imply that $-2\leq F\leq 2$ on $X.$ It only remains to slightly modify $F$ so that the final function $h_{S}$ takes values in $[0,1].$
To do so, we pick a bump $\theta:\mathbb{R}\to[0,1]$ of class $C^{\infty}(\mathbb{R})\cap C^{1,\alpha}(\mathbb{R})\cap\operatorname{Lip}(\mathbb{R})$ with $\theta(t)=0$ for $t\leq 0$ and $t\geq 2,$ and $\theta(1)=1.$ Define $h_{S}:=\theta\circ F:X\to[0,1].$ It is immediate that $h_{S}(x)=0$ for every $x\in S$, $h_{S}(x)=1$ if $d(x,S)\geq 1,$ and that $Dh_{S}=0$ on $E.$ Since $F$ is $\kappa_{1}$-Lipschitz, $h_{S}$ is $\operatorname{Lip}(\theta)\kappa_{1}$-Lipschitz. Finally, using the facts that $\operatorname{Lip}(F)+\|DF\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa_{1}$ and that $\operatorname{Lip}(\theta)+\|\theta^{\prime}\|_{\dot{C}^{0,\alpha}(\mathbb{R})}\leq c(\alpha)$ for some $c(\alpha)>0,$ it is an easy exercise to verify that
$\|Dh_{S}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\|\theta^{\prime}\|_{\dot{C}^{0,\alpha}(\mathbb{R})}\operatorname{Lip}(F)^{\alpha}+\operatorname{Lip}(\theta)\|DF\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa(\alpha,\kappa_{1}).$
Therefore, we can find $\kappa=\kappa(\alpha,C)$ for which $h_{S}$ satisfies all the properties stated in (1).
Let $g:X\to\mathbb{R}$ be $L$-Lipschitz, and $\varepsilon>0.$ We will now apply part (1) for a suitable sequence of sets to obtain the desired approximation. In order to do so, we use the same construction as in [HajJoh10, Proposition 1], replacing the $C^{1}$-separating functions from there with those we just obtained in part (1).
Define $\tilde{g}(x)=\varepsilon^{-1}g(\varepsilon x/L),$ for all $x\in X,$ and note that $\tilde{g}$ is $1$-Lipschitz. Now, we define the set $S_{n}=\{x\in X\,:\,\tilde{g}(x)\geq n\},$ for each integer $n\in\mathbb{Z}.$ Note that $S_{n+1}\subset S_{n}$, and also, because $\tilde{g}$ is $1$-Lipschitz, one has that $d(X\setminus S_{n},S_{n+1})\geq 1$ for every $n\in\mathbb{Z}.$ Now, for each subset $S_{n},$ let $h_{S_{n}}:X\to[0,1]$ be the function from (1). Define now $h_{n}=1-h_{S_{n+1}}$ for every $n$, and also
$h(x)=\sum_{n=0}^{\infty}h_{n}(x)+\sum_{n=-\infty}^{-1}(h_{n}(x)-1),\quad x\in X.$
Each $h_{n}$ is $C^{1}$ and $\kappa$-Lipschitz, $h_{n}=1$ on $S_{n+1},$ and $h_{n}(x)=0$ if $d(x,S_{n+1})\geq 1.$ As proven in [HajJoh10, Proposition 1], $h:X\to\mathbb{R}$ is a well-defined $\kappa$-Lipschitz and $C^{1}$ function, with $\sup_{X}|h-\tilde{g}|\leq 1.$ Moreover, the sums defining $h$ are locally finite, meaning that for every $x\in X,$ there exists $N_{x}\in\mathbb{N},$ and a ball $B(x,r_{x})$ so that
$h(z)=\sum_{n=0}^{\max\{|m|,N_{x}\}}h_{n}(z)+\sum_{n=-\max\{|m|,N_{x}\}}^{-1}(h_{n}(z)-1),\quad z\in B(x,r_{x}),\,m\in\mathbb{Z}.$ (37)
The relation (37) will allow us to differentiate the sums defining $h$ term by term locally around every $x\in X.$
Let us now prove that $h\in C^{1,\alpha}(X)$ with $\|Dh\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq 2\kappa.$ Recalling that by definition $Dh_{S_{n+1}}(x)=0$ whenever $x\in S_{n+1}\cup\{x\in X\,:\,d(x,S_{n+1})\geq 1\}$, it is immediate from the definition $h_{n}:=1-h_{S_{n+1}}$ that
$Dh_{n}=0\quad\text{on}\quad S_{n+1}\cup\{x\in X\,:\,d(x,S_{n+1})\geq 1\},\quad n\in\mathbb{Z}.$
Note that this implies $Dh_{n}=0$ on $X\setminus S_{n},$ as $d(X\setminus S_{n},S_{n+1})\geq 1.$ These observations together with the continuity of $Dh_{n}$ give
$Dh_{n}=0\quad\text{on}\quad S_{n+1}\cup\overline{X\setminus S_{n}},\quad n\in\mathbb{Z}.$ (38)
Also, we claim that
$Dh(x)=Dh_{m}(x),\quad\text{whenever}\quad x\in S_{m}\setminus S_{m+1},\quad m\in\mathbb{Z}.$ (39)
Indeed, let $x\in S_{m}\setminus S_{m+1}$. For those $n\geq m+1,$ we have $x\notin S_{n}$ and thus $Dh_{n}(x)=0,$ e.g., by virtue of (38). And for those $n\leq m-1,$ we have $x\in S_{n+1}$ and so $Dh_{n}(x)=0.$ In other words, $Dh_{n}(x)=0$ for $n\neq m.$ It then follows by (37) and evaluating at $x\in S_{m}\setminus S_{m+1}$ that
$Dh(x)=\sum_{n=0}^{\max\{|m|,N_{x}\}}Dh_{n}(x)+\sum_{n=-\max\{|m|,N_{x}\}}^{-1}Dh_{n}(x)=Dh_{m}(x).$
This proves the claim (39).
Now given $x,y\in X,$ let $m\in\mathbb{Z}$ and $l\in\mathbb{N}\cup\{0\}$ be so that $x\in S_{m}\setminus S_{m+1}$ and $y\in S_{m+l}\setminus S_{m+l+1},$ and we show the $\alpha$-Hölder estimate for $\|Dh(x)-Dh(y)\|_{*}.$
In the case $l=0,$ we have $x,y\in S_{m}\setminus S_{m+1},$ and so it suffices to apply (39) and the fact that $\|Dh_{m}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa.$
Assume now $l\geq 1$. By the connectedness of $[x,y]$ and $x\not\in S_{m+l}$ and $y\in S_{m+l}$, the segment $[x,y]$ must intersect the boundary of $S_{m+l}.$ Let $z\in[x,y]\cap S_{m+l}\cap\overline{X\setminus S_{m+l}}$. In particular, $z\in\overline{X\setminus S_{m+l}},$ and by (38), this implies $Dh_{m+l}(z)=0.$ Also, observe that $y\in S_{m+l}\subset S_{m+1},$ and so $Dh_{m}(y)=0,$ again by (38). Using first (39), then $Dh_{m}(y)=Dh_{m+l}(z)=0$, and finally that $\|Dh_{n}\|_{\dot{C}^{0,\alpha}(X,X^{*})}\leq\kappa$ for all $n\in\mathbb{Z},$ we conclude that
$\|Dh(x)-Dh(y)\|_{*}$ $=\|Dh_{m}(x)-Dh_{m+l}(y)\|_{*}$
$\leq\|Dh_{m}(x)-Dh_{m}(y)\|_{*}+\|Dh_{m+l}(z)-Dh_{m+l}(y)\|_{*}$
$\leq\kappa\left(\|x-y\|^{\alpha}+\|z-y\|^{\alpha}\right)\leq 2\kappa\|x-y\|^{\alpha}.$
By symmetry, we can swap the roles of $x$ and $y,$ and thus we have shown that $h\in C^{1,\alpha}(X).$
Finally, rescaling the function $h$ by $\tilde{h}(x)=\varepsilon h(xL/\varepsilon)$ and noting that $g(x)=\varepsilon\tilde{g}(xL/\varepsilon),$ we see that $\tilde{h}\in C^{1,\alpha}(X)$ with $\sup_{X}|\tilde{h}-g|\leq\varepsilon$ and $\operatorname{Lip}(\tilde{h})\leq L\operatorname{Lip}(h)\leq\kappa L.$
We will use the following elementary fact.
Remark 3.9.
If $U,V$ are normed spaces, and $\psi:U\to V$ is Lipschitz and bounded, then $\psi$ is $\alpha$-Hölder for every $\alpha\in(0,1].$
A consequence of Theorem 3.8, in combination with our results from previous sections, is the following $C^{1,\alpha}\cap\operatorname{Lip}$ approximation of $\dot{\operatorname{VC}}^{0,\omega}$ functions in super-reflexive spaces. Note that, unlike in Theorem 3.5, $X$ may be nonseparable.
Theorem 3.10 (Approximation in super-reflexive spaces).
[…]
  • (i)
    $\overline{C^{1,\alpha}(X)\cap\mathrm{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$
  • (ii)
    $\overline{C^{1,\alpha}_{\operatorname{bs}}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}(X).$
Notice that the modulus $\omega$ is independent of $\alpha.$ Also, the approximations in part (ii) are Lipschitz. Indeed, as the derivative is $\alpha$-Hölder and boundedly supported, it is in particular bounded and this implies that the function is Lipschitz.
Proof.
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\varepsilon>0.$ Applying Lemma 3.1, we can find a function $g\in\operatorname{Lip}(X)$ so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ Now, by Theorem 3.8(2), there is a sequence $(g_{n})$ of $\operatorname{Lip}(X)\cap C^{1,\alpha}(X)$ functions with $\sup_{n}\operatorname{Lip}(g_{n})\leq\kappa\operatorname{Lip}(g),$ and $\lim_{n}\sup_{X}|g-g_{n}|=0.$ Therefore Lemma 3.2 yields that $\lim_{n}\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}=0,$ and taking some $g_{n}$ from the sequence with $\|g_{n}-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ we conclude the inclusion ``""```$\supset$"` ` ⊃ " of $(i).$ The reverse inclusion follows by noting that $\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)\subset\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$ by the condition (3), and that $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ is a closed subspace of $\dot{C}^{0,\omega}(X).$
Let $f\in\dot{\operatorname{VC}}^{0,\omega}(X)$ and $\varepsilon>0.$ By Theorem 3.3 we find $g\in\operatorname{Lip}(X)$ with bounded support so that $\|f-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon.$ With $R>0$ so that $g(x)=0$ for $\|x\|\geq R,$ we use Theorem 3.8(1) to obtain a $\operatorname{Lip}(X)\cap C^{1,\alpha}(X)$ function $\varphi:X\to[0,1]$ with $\varphi=1$ on $B(0,2R)$ and $\varphi=0$ on $X\setminus B(0,2R+1).$ If $(g_{n})_{n}$ is the approximating function for $g$ from part (i), we define $h_{n}=\varphi g_{n}.$ Obviously each $h_{n}\in C^{1}(X)$ and also has bounded support, and moreover the properties $\|h_{n}-g\|_{\infty}\to 0$ and $\limsup_{n}\operatorname{Lip}(h_{n})<\infty$ hold and are checked exactly as in the proof of Lemma 3.4, see the line (34). So, it only remains to verify that $h_{n}$ has $\alpha$-Hölder derivative, at least, for $n$ large enough. Write,
$Dh_{n}(x)=g_{n}(x)D\varphi(x)+\varphi(x)Dg_{n}(x)$
and let us prove that both terms define an $\alpha$-Hölder function.
For the first function $x\mapsto g_{n}(x)D\varphi(x)$, let us observe that, for $n$ large enough, one has $\sup_{X}|g_{n}|\leq 1+\sup_{X}|g|,$ as $g_{n}$ converges uniformly on $X$ to $g.$ Hence, by Remark 3.9, $g_{n}$ is $\alpha$-Hölder continuous on $X.$ Then, we write
$\|g_{n}(x)D\varphi(x)$ $-g_{n}(z)D\varphi(z)\|_{*}$
$\leq|g_{n}(x)|\|D\varphi(x)-D\varphi(z)\|_{*}+|g_{n}(x)-g_{n}(z)|\|D\varphi(z)\|_{*}$
$\leq\sup_{X}|g_{n}|\|D\varphi\|_{\dot{C}^{0,\alpha}(X,X^{*})}\|x-z\|^{\alpha}+\|g_{n}\|_{\dot{C}^{0,\alpha}(X,\mathbb{R})}\|x-z\|^{\alpha}\sup_{X}\|D\varphi\|_{*}.$
All the factors that multiply the term $\|x-z\|^{\alpha}$ are finite by the properties of $\varphi$ and $g_{n}.$
As concerns the second function $x\mapsto\varphi(x)Dg_{n}(x)$, its $\alpha$-Hölder continuity is verified in a very similar way, this time using Remark 3.9 for $\varphi$ and that $Dg_{n}$ is bounded (as $g_{n}$ is Lipschitz) and $\alpha$-Hölder continuous.
We conclude that $(h_{n})_{n}\to g$ uniformly, with $(h_{n})_{n}\in C^{1,\alpha}_{\operatorname{bs}}(X)$ and $\limsup_{n}\operatorname{Lip}(h_{n})<\infty.$ By Lemma 3.2, we have $\|h_{n}-g\|_{\dot{C}^{0,\omega}(X)}\to 0$ as well. Therefore, we can find $h\in C^{1,\alpha}_{\operatorname{bs}}(X)$ with $\|h-g\|_{\dot{C}^{0,\omega}(X)}\leq\varepsilon,$ and thus $\|h-f\|_{\dot{C}^{0,\omega}(X)}\leq 2\varepsilon.$
We have shown the inclusion $\dot{\operatorname{VC}}^{0,\omega}(X)\subset\overline{C^{1,\alpha}_{\operatorname{bs}}(X)}^{\dot{C}^{0,\omega}(X)}.$ For the reverse inclusion, observe that if $f\in C^{1,\alpha}_{\operatorname{bs}}(X),$ then $Df:X\to X^{*}$ is $\alpha$-Hölder and with bounded support, and so $Df$ is bounded in $X.$ Because Lipschitz functions with bounded support are contained in $\dot{\operatorname{VC}}^{0,\omega}(X),$ which is closed under limits with respect to the $\dot{C}^{0,\omega}(X)$ seminorm, the reverse inclusion holds. ∎
Remark 3.11.
We exhibit some examples of classical spaces that are covered by Theorem 3.10.
  1. (1)
    If $X=L^{p},$ for $1<p<\infty,$ then (35) holds for $\alpha_{p}=p-1$ when $p\leq 2$ and for $\alpha_{p}=1$ when $p\geq 2$; for a proof see [DevilleGodefroyZizler, Corollary V.1.2]. By Theorem 3.10, this gives rise to approximations of $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ and $\dot{\operatorname{VC}}^{0,\omega}(X)$ by Lipschitz functions of class $C^{1,\alpha_{p}}(X)$.
  2. (2)
    If $X$ is a Hilbert space (separable or not), then (35) holds as an identity for $\alpha=1$ and $C=2$ (the parallelogram law), and so Theorem 3.10 gives $C^{1,1}$ approximations:
    • (i)
      $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)=\overline{C^{1,1}(X)\cap\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)},$
    • (ii)
      $\dot{\operatorname{VC}}^{0,\omega}(X)=\overline{C^{1,1}_{\operatorname{bs}}(X)}^{\dot{C}^{0,\omega}(X)}.$
3.3. Approximation of Banach-valued mappings from $c_{0}$
For an arbitrary set of indices $\mathcal{A},$ the space $c_{0}(\mathcal{A})$ consists of those elements $x=(x_{\alpha})_{\alpha\in\mathcal{A}}\in\mathbb{R}^{\mathcal{A}}$ such that for every $\varepsilon>0,$ there exists a finite subset $S=S_{x}\subset\mathcal{A}$ so that $|x_{\alpha}|\leq\varepsilon,$ whenever $\alpha\in\mathcal{A}\setminus S.$ We equip $c_{0}(\mathcal{A})$ with the norm $\|(x_{\alpha})_{\alpha\in\mathcal{A}}\|_{\infty}:=\sup_{\alpha\in\mathcal{A}}|x_{\alpha}|$ and this results in a Banach space.
Given $x=(x_{\alpha})_{\alpha\in\mathcal{A}}$ and $\alpha\in\mathcal{A},$ we denote the projection $(x)_{\beta}=((x_{\alpha})_{\alpha\in\mathcal{A}})_{\beta}:=x_{\beta}\in\mathbb{R};$ and for every finite subset $S\subset\mathcal{A},$ we denote
$P_{S}:c_{0}(\mathcal{A})\to c_{00}(\mathcal{A}),\qquad P_{S}(x)=\sum_{\alpha\in S}x_{\alpha}e_{\alpha},$
where $e_{\alpha}$ is the element of $c_{0}(\mathcal{A})$ that satisfies $(e_{\alpha})_{\beta}=\delta_{\alpha,\beta},$ for every $\beta\in\mathcal{A}.$
Theorem 3.12 (Approximation in $c_{0}(\mathcal{A})$).
For an arbitrary set of indices $\mathcal{A},$ let $X=c_{0}(\mathcal{A}),$ and let $Y$ be a Banach space. Then, the following hold:
  • (i)
    $\overline{C^{\infty}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y),$
  • (ii)
    $\overline{C^{\infty}_{\operatorname{bs}}(X,Y)\cap\dot{\operatorname{VC}}^{0,\omega}(X,Y)}^{\dot{C}^{0,\omega}(X,Y)}=\dot{\operatorname{VC}}^{0,\omega}(X,Y).$
Moreover, in the particular case $Y=\mathbb{R},$ the approximation can be taken to be Lipschitz:
  • (i)
    $\overline{C^{\infty}(X)\cap\operatorname{Lip}(X)\cap\dot{C}^{0,\omega}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X),$
  • (ii)
    $\overline{C^{\infty}_{\operatorname{bs}}(X)\cap\operatorname{Lip}(X)}^{\dot{C}^{0,\omega}(X)}=\dot{\operatorname{VC}}^{0,\omega}(X).$
One step in our proof of Theorem 3.12 will rely on the construction by Hájek and Johanis [HajJoh09, Theorem 1] (stated for Lipschitz functions), to obtain the $C^{\infty}$ approximation on $c_{0}(\mathcal{A}).$ However, the verification $\dot{C}^{0,\omega}$ convergence after our $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ condition will require more work.
Following [HajJoh09] we first construct an approximation that locally depends only on a finite number of coordinates.
Lemma 3.13.
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $\varepsilon>0.$ Then, there exists $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $r>0$ so that $\|g-f\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon,$ and for every $x\in X,$ there exists a finite subset $S=S_{x}$ of $\mathcal{A}$ for which $g(z)=g(P_{S}(z)),$ whenever $z\in B(x,r).$
Moreover, if $f$ has bounded support (resp. Lipschitz), then the above $g$ has bounded supported (resp. Lipschitz) too.
Proof.
Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ and $\varepsilon>0,$ let $\delta>0$ be so that
$\sup_{\|u-v\|\leq\delta}\frac{\|f(u)-f(v)\|_{Y}}{\omega(\|u-v\|)}\leq\frac{\varepsilon}{2}.$ (40)
By $\omega(0)=0,$ we find $r>0$ so that $2\omega(r)\|f\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon\omega(\delta).$ Let us define
$\varphi:\mathbb{R}\to\mathbb{R},\qquad\varphi(t)=\left\{\begin{array}[]{ccl}t+r&\mbox{if }&t\leq-r,\\ 0&\mbox{if }&-r\leq t\leq r,\\ t-r&\mbox{if }&t\geq r,\end{array}\right.$
and
$\phi:X\to X,\qquad\phi(x)=\phi((x_{\alpha})_{\alpha\in\mathcal{A}})=(\varphi(x_{\alpha}))_{\alpha\in\mathcal{A}}.$
As clearly $\varphi$ is $1$-Lipschitz, it is immediate that $\phi$ is $1$-Lipschitz. Also, it is immediate that $\|f-f\circ\phi\|_{Y}\leq\omega(r)\|f\|_{\dot{C}^{0,\omega}(X,Y)}.$ Defining $g:=f\circ\phi,$ we have $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y),$ by virtue of Lemma 2.3. If $x,z\in X$ are such that $\|x-z\|\leq\delta,$ then $\|\phi(x)-\phi(z)\|\leq\delta$ and so
$\|g(x)-g(z)\|_{Y}=\|f(\phi(x))-f(\phi(z))\|_{Y}\leq\frac{\varepsilon}{2}\omega(\|\phi(x)-\phi(z)\|)\leq\frac{\varepsilon}{2}\omega(\|x-z\|).$
This estimate, in combination with (40), yields,
$\sup_{\|x-z\|\leq\delta}\frac{\|(f-g)(x)-(f-g)(z)\|_{Y}}{\omega(\|x-z\|)}\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon.$
On the other hand, the choice of $r$ permits to estimate by
$\sup_{\|x-z\|\geq\delta}\frac{\|(f-g)(x)-(f-g)(z)\|_{Y}}{\omega(\|x-z\|)}\leq\frac{2\|f-g\|_{\infty}}{\omega(\delta)}\leq\frac{2\omega(r)\|f\|_{\dot{C}^{0,\omega}(X,Y)}}{\omega(\delta)}\leq\varepsilon.$
We have shown that $\|f-g\|_{\dot{C}^{0,\omega}(X,Y)}\leq\varepsilon.$
Now, if $x\in X=c_{0}(\mathcal{A}),$ then (by definition) there exists a finite subset $S=S_{x}$ such that if $\alpha\in\mathcal{A}\setminus S,$ then $|x_{\alpha}|\leq r/2.$ Then, for $z\in B(x,r/2)$ and for each $\alpha\in\mathcal{A}\setminus S$ it follows that $|z_{\alpha}|\leq r,$ implying that $\varphi(z_{\alpha})=0.$ It follows for $z\in B(x,r/2)$ that $\phi(z)=\phi(P_{S}(z))$ and hence that $g(z)=g(P_{S}(z)).$
For the second part, suppose there exists $R>0$ so that $f(x)=0$ for all $\|x\|\geq R.$ If $\|z\|\geq R+r,$ then $|z_{\alpha_{0}}|\geq R+r$ for some $\alpha_{0}\in\mathcal{A}.$ Consequently
$\|\phi(z)\|=\sup_{\alpha\in\mathcal{A}}|\varphi(z_{\alpha})|\geq|\varphi(z_{\alpha_{0}})|\geq|z_{\alpha_{0}}|-r\geq R,$
and thus $g(z)=f(\phi(z))=0.$ Finally, if $f$ is Lipschitz, then clearly $g=f\circ\phi$ is Lipschitz.
Proof of Theorem 3.12.
In the proof of part (i), by Lemma 3.13, given a fixed $r>0,$ it is enough to approximate functions $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with the property that for every $x\in X,$ there exists finite $S_{x}\subset\mathcal{A}$ such that $g(z)=g(P_{S_{x}}(z)),$ for all $z\in B(x,2r).$
Given $\varepsilon,$ let $\delta>0$ be so that
$\|g(u)-g(v)\|_{Y}\leq\varepsilon\omega(\|u-v\|),\quad\text{whenever}\quad\|u-v\|\leq\delta.$ (41)
Moreover, for $0<\eta<\min(r,\varepsilon)$ small enough (more precisely, so that $\omega(\eta)\leq\varepsilon\omega(\delta)/(1+\|g\|_{\dot{C}^{0,\omega}})),$ let $\theta:\mathbb{R}\to\mathbb{R}$ be an even $C^{\infty}$ smooth non-negative function such that $\int_{\mathbb{R}}\theta=1$ and $\theta=0$ on $\mathbb{R}\setminus[-\eta,\eta].$ For every finite set $F\subset\mathcal{A},$ and $x\in X,$ we define the Bochner integral
$h_{F}(x):=\int_{[-\eta,\eta]^{|F|}}g\Big{(}x-\sum_{\alpha\in F}t_{\alpha}e_{\alpha}\Big{)}\prod_{\alpha\in F}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|F|}(t),$ (42)
where $\lambda_{|F|}$ is the Lebesgue measure on $\mathbb{R}^{|F|}.$
Because $g(z)=g(P_{S_{x}}(z))$ for all $z\in B(x,2r)$, the fact that $\eta<r$ implies that $h_{S_{x}}=h_{S_{x}}\circ P_{S_{x}}$ on the ball $B(x,r).$ Therefore
$h_{S_{x}}(z)=\int_{[-\eta,\eta]^{|S_{x}|}}g\Big{(}P_{S_{x}}(z)-\sum_{\alpha\in S_{x}}t_{\alpha}e_{\alpha}\Big{)}\prod_{\alpha\in S_{x}}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S_{x}|}(t),\quad z\in B(x,r).$
Since $g\in\dot{C}^{0,\omega}(X,Y)$ and $P_{S_{x}}$ is $1$-Lipschitz, the function $z\mapsto(g\circ P_{S_{x}})(z)$ belongs to $\dot{C}^{0,\omega}(X,Y)$, implying that $z\mapsto(g\circ P_{S_{x}})(z)$ is uniformly continuous on $X.$ Therefore, by the previous formula, $B(x,r)\ni z\mapsto h_{S_{x}}(z)$ is a finite dimensional smooth mollification (in the Bochner sense) with a uniformly continuous mapping, which shows that $h_{S_{x}}\in C^{\infty}(B(x,r),Y).$ Moreover, using Fubini’s theorem, one can easily check that if $F\subset\mathcal{A}$ is finite and $F\supset S_{x},$ then also $h_{F}=h_{S_{x}}$ on $B(x,r).$ This enables us to define the desired mapping $h:X\to Y$ in the following manner. Given $x\in X,$ let $S_{x}\subset\mathcal{A}$ be the subset described above, and define $h(x):=h_{S_{x}}(x).$ Moreover, ordering the collection of finite subsets $\mathcal{F}(\mathcal{A})$ of $\mathcal{A}$ by inclusion, the pointwise limit of the net $\{h_{S}\}_{S\in\mathcal{F}(\mathcal{A})}$ is exactly $h.$
Now, for every fixed ball $B(x,r)$ and $z\in B(x,r)$, let $F=S_{x}\cup S_{z}$. By the definition of $h$ and the above mentioned properties, we derive
$h(z)=h_{S_{z}}(z)=h_{F}(z)=h_{S_{x}}(z).$
Since $z\mapsto h_{S_{x}}(z)$ is of class $C^{\infty}(B(x,r),Y)$, the above yields $h\in C^{\infty}(X,Y).$ All these properties were stated and proved in [HajJoh09, Lemma 6].
We next show that $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Given any two points $x,z\in X,$ let $S_{x},S_{z}\subset\mathcal{A}$ be the finite subsets associated with $x$ and $z,$ and set $S=S_{x}\cup S_{z}.$ Since $h(x)=h_{S}(x)$ and $h(z)=h_{S}(z)$ we have
$\begin{split}\|h(x)-h(z)\|_{Y}&=\Big{\|}\int_{[-\eta,\eta]^{|S|}}\Big{(}g\big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}-g\big{(}z-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}\Big{)}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)\Big{\|}_{Y}\\ &\leq\int_{[-\eta,\eta]^{|S|}}\Big{\|}g\big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}-g\big{(}z-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\big{)}\Big{\|}_{Y}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)\\ &\leq\sup_{\begin{subarray}{c}u,v\in X,\\ \|u-v\|=\|x-z\|\end{subarray}}\|g(u)-g(v)\|_{Y},\end{split}$ (43)
bearing in mind that $\int_{[-\eta,\eta]^{|S|}}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)=1.$ It follows immediately from the bound (43) that $\|h\|_{\dot{C}^{0,\omega}(X,Y)}\leq\|g\|_{\dot{C}^{0,\omega}(X,Y)}$ and $h\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X,Y).$ Now if $\|x-z\|\leq\delta,$ then together (43) and (41) give
$\|(g-h)(x)-(g-h)(z)\|_{Y}\leq\|g(x)-g(z)\|_{Y}+\|h(x)-h(z)\|_{Y}\leq 2\varepsilon\omega(\|x-z\|).$
Hence to verify $\|g-h\|_{\dot{C}^{0,\omega}(X,Y)}\leq 2\varepsilon,$ it remains to check the case $\|x-z\|\geq\delta.$ To do so, let us see that $\sup_{X}\|g-h\|\leq\varepsilon\omega(\delta).$ Indeed, if $x\in X$ and $S=S_{x},$ then again $h(x)=h_{S_{x}}(x),$ and we obtain
$\|h(x)-g(x)\|_{Y}$ $\leq\int_{[-\eta,\eta]^{|S|}}\Big{\|}g\Big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{)}-g(x)\Big{\|}_{Y}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)$
$\leq\|g\|_{\dot{C}^{0,\omega}(X,Y)}\int_{[-\eta,\eta]^{|S|}}\omega\Big{(}\Big{\|}\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{\|}\Big{)}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)$
$\leq\|g\|_{\dot{C}^{0,\omega}(X,Y)}\omega(\eta)\leq\varepsilon\omega(\delta),$
where we chose $\eta$ small enough in the beginning of the proof for the last bound to hold. Now, if $x,z\in X$ are such that $\|x-z\|\geq\delta,$ then
$\|(g-h)(x)-(g-h)(z)\|\leq 2\sup_{X}\|g-h\|\leq 2\varepsilon\omega(\delta)\leq 2\varepsilon\omega(\|x-z\|).$
We conclude that $\|g-h\|_{\dot{C}^{0,\omega}(X,Y)}\leq 2\varepsilon$ and complete the proof of Theorem 3.12(i).
For part (ii), given $f\in\dot{\operatorname{VC}}^{0,\omega}(X,Y),$ by Theorem 1.2, we approximate $f$, in the $\dot{C}^{0,\omega}(X,Y)$-seminorm by a $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X,Y)$ with bounded support. By Lemma 3.13, we can assume that for every $x\in X$ there exists a finite $S=S_{x}$ such that $g=g\circ P_{S}$ on $B(x,r),$ and for a uniform fixed $r>0.$ Let us next check that the approximation $h$ from part (i) has bounded support. Let $R>0$ be so that $g$ is zero on $X\setminus B(0,R).$ Assume that $\|x\|\geq R+\eta.$ There exists a finite set $S=S_{x}\subset\mathcal{A}$ so that $h(x)=h_{S}(x)$. Now, for $t=(t_{\alpha})_{\alpha\in S}\in[-\eta,\eta]^{|S|},$ there holds that
$\Big{\|}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{\|}\geq\|x\|-\Big{\|}\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{\|}\geq\|x\|-\eta\geq R$
and thus
$h(x)=h_{S}(x)=\int_{[-\eta,\eta]^{|S|}}g\Big{(}x-\sum_{\alpha\in S}t_{\alpha}e_{\alpha}\Big{)}\prod_{\alpha\in S}\theta(t_{\alpha})\,\mathrm{d}\lambda_{|S|}(t)=0.$
The proof of Theorem 3.12(ii) is now complete.
Finally, in the particular case $Y=\mathbb{R},$ let us establish the desired Lipschitz approximations. Let $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(X)$ be the function to be approximated. By Theorem 3.3 we can assume that $f$ is Lipschitz, and by Lemma 3.13, there is a Lipschitz approximation $g$ of $f$, with $g$ locally depending only on a finite number of coordinates around each ball of fixed radius $r>0.$ Repeating the construction of $h\in C^{\infty}(X,)$, we immediately see from the estimate (43), that $h$ is Lipschitz, and in fact $\operatorname{Lip}(h)\leq\operatorname{Lip}(g).$ Because any Lipschitz mapping in $\dot{C}^{0,\omega}(X)$ belongs to $\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(X),$ by the condition (3), the identity in $(i)$ now follows at once.
As for approximation of an $f\in\dot{\operatorname{VC}}^{0,\omega}(X)$, we apply Theorem 3.3 to reduce matters to $f\in\operatorname{Lip}_{\operatorname{bs}}(X)$. Then the function $g$ of Lemma 3.13 can be taken Lipschitz and with bounded support, and so can the function $h$ we constructed in (i) and (ii) of the present theorem. Thus $h\in C^{\infty}_{\operatorname{bs}}(X)\cap\operatorname{Lip}(X)$ approximates $f$ in the $\dot{C}^{0,\omega}(X)$ seminorm. Because any Lipschitz mapping with bounded support belongs to $\dot{\operatorname{VC}}^{0,\omega}(X),$ the identity in (ii) now follows at once.
For the particular case of $\mathcal{A}=\{1,\ldots,n\},$ $c_{0}(\mathcal{A})$ becomes $\mathbb{R}^{n}$ with the supremum norm, which is of course equivalent with the usual Euclidean setting. Theorem 3.12 then has the following consequence.
Corollary 3.14.
Let $Y$ be an arbitrary Banach space. Then:
  • (i)
    $\overline{C^{\infty}(\mathbb{R}^{n},Y)\cap\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}=\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}(\mathbb{R}^{n},Y),$
  • (ii)
    $\overline{C^{\infty}_{c}(\mathbb{R}^{n},Y)}^{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}=\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$
Appendix A Comparison of $\dot{C}^{0,\omega}$ and uniform convergence
We begin with a remark concerning the two types of convergences we are dealing with. We will refer to the norm $|||\cdot|||_{C^{0,\omega}(X,Y)}$ defined in (1).
Remark A.1.
Let $X$ and $Y$ be normed spaces and $\omega$ any modulus of continuity. Suppose that a sequence of functions $(f_{n})_{n}\subset\dot{C}^{0,\omega}(X,Y)$ converges to $f,$ with respect to the norm $|||\cdot|||_{C^{0,\omega}(X,Y)}.$ Then also $f_{n}\to f$ uniformly on bounded subsets of $X.$
Indeed, let $B$ a bounded subset of $X.$ Given $\varepsilon>0,$ let $N$ be so that $|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\varepsilon/(1+\omega(\operatorname{diam}(B)))$ for all $n\geq N.$ Then, for those $n\geq N,$ and $x\in B,$ we can write
$|f(x)-f_{n}(x)|$ $\leq|f(0)-f_{n}(0)|+|(f-f_{n})(x)-(f-f_{n})(0)|$
$\leq\left(1+\omega(\|x\|)\right)|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\left(1+\omega(\operatorname{diam}(B))\right)|||f_{n}-f|||_{C^{0,\omega}(X,Y)}\leq\varepsilon.$
However, the convergence in $\dot{C}^{0,\omega}$ does not imply global uniform convergence, even if we assume that the pertinent sequence of functions is globally uniformly bounded, as the next example shows.
Example A.2.
Let $\omega$ be of the form $\omega(t)=t^{\alpha},$ with $0<\alpha\leq 1$ Define $f_{n}:\mathbb{R}\to\mathbb{R}$ as the continuous function given by $f_{n}=0$ on $(-\infty,0]\cup[2n,+\infty),$ $f_{n}(n)=1$, and $f_{n}$ affine on both $[0,n]$ and $[n,2n].$ It is immediate that $\|f_{n}\|_{\infty}\leq 1$ for all $n,$ and that $(f_{n})_{n}$ does not converge to 00 uniformly on $\mathbb{R}.$ However, one has $f_{n}\to 0$ in the $\dot{C}^{0,\omega}(\mathbb{R})$ sense. This easily follows after observing that
$\sup_{n\leq s<t\leq 2n}\frac{|f_{n}(t)-f_{n}(s)|}{\omega(|t-s|)}=\sup_{0\leq s<t\leq n}\frac{|f_{n}(t)-f_{n}(s)|}{\omega(|t-s|)}=\sup_{0\leq s<t\leq n}\frac{1}{n}\frac{t-s}{(t-s)^{\alpha}}=\frac{1}{n^{\alpha}}.$
Finally, we next show that the converse to Remark A.1 is false, by providing a sequence $(f_{n})_{n}\subset\dot{C}^{0,\omega}$ and $f\in\dot{C}^{0,\omega}$ so that $f_{n}$ converges uniformly on $\mathbb{R},$ and $f_{n}$ does not converge to $f$ in the $\|\cdot\|_{\dot{C}^{0,\omega}}$ norm on bouded subsets.
Example A.3.
Define $f=0$ on $\mathbb{R}$ and the continuous functions $f_{n}:\mathbb{R}\to[0,1]$ given by
$f_{n}(t)=\begin{cases}1/\sqrt{n},\qquad&t=0,\\ \text{affine},\qquad&t\in[0,1/n],\\ 0,\qquad&t\geq 1/n,\\ f_{n}(-t),\qquad&t\leq 0.\end{cases}$
Then $\sup_{t\in\mathbb{R}}|f_{n}(t)|=1/\sqrt{n},$ yielding $\|f_{n}-f\|_{\infty}\to 0.$ Also, each $f_{n}$ is $\sqrt{n}$-Lipschitz:
$\frac{|f_{n}(0)-f_{n}(1/n)|}{|0-1/n|}=\frac{1/\sqrt{n}}{1/n}=\sqrt{n}.$
Because $f_{n}$ is bounded, then $f_{n}\in\dot{C}^{0,\omega}(\mathbb{R})$ for all moduli $\omega$ that satisfy $\lim_{t\to 0}t/\omega(t)<\infty.$
However, if $\omega(t)=t^{\alpha}$ with $\alpha\in(1/2,1],$ then $\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}([0,1])}=\infty,$ since
$\lim_{n}\|f_{n}-f\|_{\dot{C}^{0,\omega}([0,1])}\geq\lim_{n}\frac{|f_{n}(0)-f_{n}(1/n)|}{\omega(|0-1/n|)}=\lim_{n}n^{\alpha-\frac{1}{2}}=\infty.$